### v2

# MultiConIR: Towards multi-condition Information Retrieval

Abstract
In this paper, we introduce MultiConIR, the first benchmark designed to evaluate retrieval models in multi-condition scenarios. Unlike existing datasets that primarily focus on single-condition queries from search engines, MultiConIR captures real-world complexity by incorporating five diverse domains: books, movies, people, medical cases, and legal documents. We propose three tasks to systematically assess retrieval and reranking models on multi-condition robustness, monotonic relevance ranking, and query format sensitivity. Our findings reveal that existing retrieval and reranking models struggle with multi-condition retrieval, with rerankers suffering severe performance degradation as query complexity increases. We further investigate the performance gap between retrieval and reranking models, exploring potential reasons for these discrepancies, and analysis the impact of different pooling strategies on condition placement sensitivity. Finally, we highlight the strengths of GritLM and Nv-Embed, which demonstrate enhanced adaptability to multi-condition queries, offering insights for future retrieval models. The code and datasets are available at https://github.com/EIT-NLP/MultiConIR
MultiConIR: Towards multi-condition Information Retrieval
Xuan Lu1,2, Sifan Liu2, Bochao Yin2, Yongqi Li3, Xinghao Chen2,3, Hui Su4, Yaohui Jin1, Wenjun Zeng2, Xiaoyu Shen2thanks: Corresponding author, 1Shanghai Jiao Tong University 2Ningbo Key Laboratory of Spatial Intelligence and Digital Derivative, Institute of Digital Twin, EIT 3Department of Computing, The Hong Kong Polytechnic University 4Meituan Inc. xyshen@eitech.edu.cn
1 Introduction
Information retrieval (IR) is critical for helping users find relevant information across various domains. Traditionally, IR systems retrieve documents by matching queries based on lexical similarity, such as BM25 (Carpineto and Romano, 2012; Ponte and Croft, 2017), or semantic similarity using dense vector representations (Karpukhin et al., 2020; Zhan et al., 2021). Though highly effective for queries with straightforward query-document relationships (Su et al., 2024), they often fail to fully capture nuanced intent as user needs become more complex (Zhu et al., 2023; Su et al., 2024).
Figure 1: From single-condition to multi-condition retrieval. Standard retrieval relies on single-condition queries, while instruction-following retrieval enhances queries with explicit instructions. Traditional SQL-based filtering handles predefined attributes, whereas real-world multi-condition retrieval involves queries specifying multiple conditions, which can be attribute-level or semantic-level.
A significant challenge arises when users specify multiple requirements simultaneously, as illustrated in Fig 1. Whether searching for a movie with specific attributes or selecting a product that meets various criteria, multi-condition search has become an integral part of modern information-seeking behavior. Traditional IR systems handle such scenarios using structured filtering, such as SQL-based queries that retrieve information from backend databases based on predefined conditions. However, this approach is inherently rigid and limited, as it relies on explicitly defined attributes and lacks the flexibility to accommodate evolving or diverse user preferences. As a result, it struggles to support nuanced or semantic-level queries that go beyond structured data filtering.
The advent of Large Language Models (LLMs) has enhanced IR by introducing instruction-following capabilities (Asai et al., 2023; Weller et al., 2024a; Oh et al., 2024). This approach augments standard queries with explicit instructions, which serve as additional constraints to refine search results, as shown in Fig.1. Despite these advancements, existing evaluation benchmarks remain predominantly focused on single-condition queries and binary relevance assessments—classifying documents as either relevant or irrelevant (Nguyen et al., 2016; Kwiatkowski et al., 2019; Muennighoff et al., 2022)—thus overlooking the nuanced challenges of multi-condition queries, where relevance depends on the degree to which multiple conditions are satisfied.
An ideal multi-condition retrieval system should exhibit the following properties: (1) Complexity Robustness: The system should maintain high performance regardless of query complexity (i.e., the number of conditions specified); (2) Relevance Monotonicity: The relevance scores should scale monotonically with the number of matched conditions; for example, a document matching all $n$ conditions should be ranked higher than one matching $n-1$; (3) Format Invariance: The system should yield consistent results regardless of the query format, whether presented as a structured list or as free-form natural language.
Existing benchmarks do not offer a structured framework for evaluating multi-condition retrieval along these dimensions. To address this gap, we introduce MultiCondIR—the first benchmark designed to comprehensively evaluate multi-condition retrieval systems. Through systematic experiments on 12 state-of-the-art models (including dense retrievers, cross-encoders, and LLM-based agents), we uncover several critical insights:
  • Multi-Condition Degradation: Dense retrieval and reranking models deteriorate as query conditions increase, with rerankers suffering more severely. Rerankers’ token-level interactions benefit single-condition tasks but hinder multi-condition retrieval.
  • Relevance Monotonicity Failure: Models fail to maintain a consistent ranking hierarchy. Dense retrievers tend to focus more on “exact match” and “complete mismatch”, likely due to their training paradigm.
  • Query Format Sensitivity: Models are susceptible to query phrasing, with rerankers exhibiting a flip rate exceeding 20%.
  • Pooling Impact on Focus: Different embedding pooling strategies influence the model’s focus on query condition positions. Mean pooling focuses on early query conditions, <EOS> pooling biases toward later conditions, while latent attention layer pooling distributes focus more evenly.
  • GritLM’s Advantage: GritLM outperforms other models, potentially due to its hybrid attention mechanism, which integrates both bidirectional and causal attention to improve the understanding of complex queries.
By quantifying these gaps, our work provides actionable guidelines for developing IR systems that truly understand multi-condition intent, laying the groundwork for advancing IR toward human-like reasoning in complex search scenarios.
2 Related Works
Retriever: From Sparse To Dense
Traditional sparse retrieval methods are based on BM25 (Robertson and Zaragoza, 2009), TF-IDF (Ramos et al., 2003), etc., rely on keyword matching and statistical weighting to evaluate relevance, which suffers from the well-known issue of lexical gap (Berger et al., 2000), restricting their ability to effectively capture semantic relationships (Luan et al., 2021; Nian et al., 2024).
Dense retrieval addresses this limitation by encoding both queries and documents as embeddings within a joint latent space, where the semantic relationship is captured through the similarity scores between their embeddings (Li et al., 2023a). Pre-trained language models like BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020), are widely used as backbone encoders for dense retrieval (Li et al., 2023b; Sturua et al., 2024; Xiao et al., 2023).
Recent advancements have shown that LLMs offer significant potential as backbone encoders for dense retrieval (Wang et al., 2024a; Weller et al., 2024c; BehnamGhader et al., 2024). For instance, Repllama (Ma et al., 2023) enhanced retrieval performance by fine-tuning Llama-2 to serve as dense retrievers. GritLM (Muennighoff et al., 2024) unified text embedding and generation within a single LLM. LLM2Vec (BehnamGhader et al., 2024) introduced an unsupervised approach for transforming decoder-only LLMs into dense retrievers.
Benchmarks In Complex Retrieval Tasks
Existing datasets for information retrieval, such as MS MARCO (Nguyen et al., 2016), Natural Questions (Kwiatkowski et al., 2019), and MTEB (Muennighoff et al., 2022), primarily focus on queries sourced from search engines. The relationships between queries and documents are typically simple and direct (Su et al., 2024).
Recent studies have expanded retrieval benchmarks to address more complex scenarios. Instruction-based datasets (Weller et al., 2024a; Qin et al., 2024; Oh et al., 2024), for instance, evaluate the instruction-following capabilities of retrieval models by embedding explicit instructions within queries to better represent users’ retrieval intents. Furthermore, some works have assessed retrieval models’ abilities to handle logical reasoning tasks, including Boolean logic Mai et al. (2024); Zhang et al. (2024c), negation Zhang et al. (2024a); Weller et al. (2024b), and multi-hop reasoning Su et al. (2024). These efforts mark significant progress in increasing query complexity. However, while research in the generative modeling domain has explored the ability of LLMs to handle multi-constraint instructions (He et al., 2024; Ferraz et al., 2024; Zhang et al., 2024b), studies on retrieval models in multi-condition scenarios remain sparse.
3 MultiConIR
We introduce MultiConIR, a benchmark designed to evaluate the capacity of retrieval models to process multi-condition queries. Formally, given a query $q_{k}$ composed of $k$ conditions $C=\{c_{1},c_{2},\dots,c_{k}\}$ with $k\in\{1,\dots,10\}$, we construct a structured retrieval setup consisting of:
(1) Two query formulations, denoted as $q_{k}^{\text{inst}}$ and $q_{k}^{\text{desc}}$, where $q_{k}^{\text{inst}}$ corresponds to a structured instruction-style query, formally expressed as a tuple $q_{k}^{\text{inst}}=\langle f,C\rangle$ where $f$ is an explicit function describing retrieval constraints, and $q_{k}^{\text{desc}}$ is a natural language descriptive-style query, represented as an unstructured sequence about the same set $C$,
(2) A positive document $d^{+}$ that satisfies all $k$ conditions, i.e., $d^{+}\models C$,
(3) A sequence of hard negative documents $\{d_{0},d_{1},\dots,d_{k-1}\}$, where each $d_{j}$ satisfies exactly $j$ out of $k$ conditions, formally expressed as $d_{j}\models\{c_{1},\dots,c_{j}\}$ and $d_{j}\not\models\{c_{j+1},\dots,c_{k}\}$.
This controlled design enables a principled evaluation of multi-condition retrieval along three fundamental axes: (1) Complexity Robustness: The model’s retrieval effectiveness as $k$ increases, measured by its ability to distinguish $d^{+}$ from $d_{k-1}$; (2) Relevance Monotonicity: The extent to which the retrieval model enforces a strict ordering such that $S(q_{k},d_{j})>S(q_{k},d_{j+1})$ for all $j$, ensuring that documents satisfying more conditions are ranked higher; and (3) Format Invariance: The stability of retrieval performance under transformations of query representation, quantified by discrepancies in ranking outcomes across query formats.
3.1 Domain Selection
MultiConIR was constructed from real-world datasets from five diverse domains: books, movies, people, medical cases, and legal documents. These domains were carefully chosen for their practical significance and their intrinsic requirement for multi-condition retrieval. Books & Movies: Queries in these domains often blend attribute-based (e.g., a sci-fi movie) with narrative-based (e.g., a story about time travel), requiring semantic understanding beyond keyword matching. People: Queries are often based on vague recollections or specific but incomplete information about a person (e.g., a Nobel laureate in Physics who studied black holes) Medical Case & Legal Document: These domains require fine-grained condition matching, where case-based reasoning is essential (e.g., a patient with breathing issues and antibiotic allergies), demanding models understand condition dependencies, not just match isolated terms.
3.2 Dataset Construction Pipeline
Figure 2: MultiCon Dataset Construction Pipeline.
To construct MultiConIR, we design a multi-step data generation framework as shown in Fig. 2, this pipeline is highly adaptable across multiple domains, enabling the generation of queries and hard negative (HN) documents that progressively satisfy 1 to 10 conditions. To preserve dataset integrity and mitigate the generalization issues associated with fully synthetic datasets (Li et al., 2023c; Wang et al., 2024b), we employ LLM-based generation (GPT-4o) exclusively for modifying sentences within hard negatives, rather than altering entire documents. [^1] Discussion about the challenges associated with fully LLM-generated datasets can be found in Appendix E.1. The data creation process consists of the following steps, with detailed prompt templates provided in Appendix B:
Step 1: Condition Sentence Extraction. To capture the fine-grained constraints for multi-condition retrieval, we prompt GPT-4o to extract ten key sentences from each document, where each sentence represents a distinct and semantically complete condition. To ensure robustness, extracted sentences must be contextually relevant, non-redundant, and representative of the document’s key information.
Step 2: Query Generation. Given the extracted condition sentences, we prompt GPT-4o to generate ten hierarchical queries (Query1 to Query10), each incorporating an increasing number of conditions (from 1 to 10). To enhance linguistic diversity, queries are formulated in two distinct styles: (1) Instruction-style (explicitly listing conditions in a structured format). (2) Descriptive-style (embedding conditions naturally within a coherent sentence).
Step 3: Hard Negative Sentence Construction. To create hard negatives, we prompt GPT-4o to modify extracted condition sentences into Hard Negative Sentences (HNS) that subtly deviate from the original meaning while preserving lexical plausibility. Unlike naive negation, these modifications introduce semantic shifts that retain fluency but fail to satisfy the original retrieval conditions, thereby making retrieval harder. We experimented with two approaches for HNS construction: one modifies key information (applied to books, movies, medical cases, and legal documents), while the other retains keywords but adds dummy information (used for the people dataset).  [^2] The differences between these two approaches are discussed in Appendix E.2.
Step 4: HN Document Generation. To construct multi-condition retrieval datasets, we systematically assemble documents: (1) Positive documents: Composed of all ten original condition sentences, ensuring full query relevance. (2) Hard Negative documents $d_{1}$ to $d_{10}$: Generated by progressively replacing 1 to 10 original condition sentences with their corresponding hard negative versions, creating a continuum of relevance degradation. This structured data generation ensures that each query condition incrementally refines document ranking, allowing for fine-grained retrieval performance assessment.
3.3 Evaluation Metrics
Complexity Robustness
Queries range from query1 to query10, each progressively incorporating 1 to 10 conditions. The candidate set comprises a Positive document that fully satisfies all conditions and a HN1 document, which is derived from the positive by modifying a single condition.
Complexity robustness is measured using Win Rate (WR) under various $k$, defined as:
$\text{WR}_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(S(q_{k},d^{+})>S(q_{k},d_{k-1})),$ (1)
where $S(q_{i},d^{+})$ and $S(q_{i},d^{-})$ denote similarity scores for the positive document and hard negative.
Relevance Monotonicity
The query is fixed as query10 (containing all 10 conditions), while the candidate set includes one positive and ten hard negatives ($d_{0}-d_{9}$), each containing 0–9 conditions.
We evaluate performance using Win Rate between adjacent hard negatives:
$\begin{split}\text{WR}_{k,k-1}&=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big{(}S(q_{10},d_{k})\\ &\quad>S(q_{10},d_{k-1})\big{)},\end{split}$ (2)
Format invariance.
We compare two query formats: (1) Instruction-style, which explicitly lists conditions (e.g., Find a movie that meets the following conditions: 1. Action genre, 2. Directed by James Cameron). (2) Descriptive-style, which integrates conditions into a natural query (e.g., Find an action movie directed by James Cameron).
To quantify ranking variability between query styles, we define the Flip Rate (FR):
$\begin{split}\text{FR}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big{(}\mathbf{1}(S_{\text{inst}}(q_{10},d_{k})>S_{\text{inst}}(q_{10},d_{k-1}))\\ \neq\mathbf{1}(S_{\text{desc}}(q_{10},d_{k})>S_{\text{desc}}(q_{10},d_{k-1}))\big{)},\end{split}$ (3)
where $S_{\text{inst}}$ and $S_{\text{desc}}$ denote similarity scores under instruction-style and descriptive-style queries. The indicator function returns 1 if the ranking order of positive and hard negative documents changes between query styles and 0 otherwise. A higher FR indicates greater sensitivity to query formulation.
Figure 3: Effect of Condition Quantity on Average Win Rate (reflecting complexity robustness). Win Rate measures Positive vs. Hard Negative success as query conditions increase. Blue, green and red lines represent Dense Retrievers, Rerankers and Red line respectively.
Figure 4: Relevance Monotonicity Distinction. Win Rate reflects the success rate between documents satisfying different numbers of conditions under a multi-condition query (query10), i.e., $d_{k}$ vs. $d_{k-1}$.
4 Experiments
We evaluate 13 representative retrieval models from diverse architectures and varying model sizes, including one spare retrieval model: BM25 (Robertson and Zaragoza, 2009); Two BERT-based retrieval models: gte-large-en-v1.5 (Li et al., 2023b) and jina-embeddings-v3 (Sturua et al., 2024); Seven LLM-based retrieval models: NV-Embed-v2 (Lee et al., 2024), bge-en-icl (Li et al., 2024), gte-Qwen2-7B-instruct (Li et al., 2023b), gte-Qwen2-1.5B-instruct (Li et al., 2023b), e5-mistral-7b-instruct (Wang et al., 2024a), GritLM-7B (Muennighoff et al., 2024), LLM2Vec (BehnamGhader et al., 2024); Three reranking models: bge-reranker-v2-m3 (Chen et al., 2024), bge-reranker-v2-gemma (Chen et al., 2024), FollowIR-7B (Weller et al., 2024a). Details of each model are provided in Appendix A.
Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 Decline
Spare Retriever
BM25 28.59 34 33.14 36.53 37.38 37.95 38.08 38.41 38.86 39.87 -11.28
Dense Retriever
jina-embeddings-v3 76.09 71.26 71.84 71.04 65.59 65.65 64.75 64.24 64.62 60.71 15.38
gte-large-en-v1.5 75.87 77.26 73.79 70.22 70.71 67.40 67.53 64.97 65.36 61.36 14.51
NV-Embed-v2 80.53 80.32 78.81 75.70 75.68 72.61 73.28 71.54 70.00 68.02 12.51
bge-en-icl 83.42 80.65 78.44 76.77 74.54 73.00 74.23 73.25 69.70 68.00 15.42
gte-Qwen2-7B-instruct 70.75 72.22 69.99 68.51 65.20 63.53 62.22 62.20 59.15 56.17 14.58
gte-Qwen2-1.5B-instruct 73.64 74.97 72.23 71.37 69.94 67.46 66.92 64.64 63.65 58.68 14.96
e5-mistral-7b-instruct 75.05 70.85 68.18 67.45 63.70 61.60 59.70 59.07 57.85 58.12 16.93
GritLM-7B 82.08 80.32 78.38 76.40 76.40 73.50 75.69 74.62 74.53 75.95 6.13
LLM2Vec 83.13 77.42 75.49 75.48 72.49 72.56 70.56 70.73 68.71 67.00 16.13
Reranker
bge-reranker-v2-m3 87.14 85.56 78.62 76.05 74.29 68.41 67.86 59.48 55.59 44.87 42.27
bge-reranker-v2-gemma 91.07 90.02 86.70 84.99 83.17 79.00 75.89 72.29 67.11 56.09 34.98
followIR 83.41 79.72 76.25 74.60 70.12 67.94 62.62 55.93 48.59 43.52 39.89
Table 1: Impact of increasing condition quantity in queries on average Win Rate (Task 1). The Decline reflects the degree of Win Rate reduction from query1 to query10.
4.1 Results for Complexity Robustness
Table 1 presents the average Win Rate scores for evaluating complexity robustness across five datasets [^3] Individual dataset performances are detailed in the Appendix D.1.. To visualize how model performance evolves as the number of query conditions increases, we plot a line graph with Gaussian smoothing (window size = 1), as shown in Fig. 4. The results reveal several notable trends:
Performance decline with increasing query conditions
As the number of conditions in the query increases, the performance of both retrieval and reranking models declines. This suggests that with more conditions, models struggle to accurately distinguish between positives and HNs. Among all models, GritLM-7B exhibits the lowest performance degradation, declining by only 6.13% from Query1 to Query10. NV-Embed-v2 follows with a relatively low decline of 12.51%.
Reranking models exhibit steeper performance drop
Initially, reranking models outperform retrieval models with single-condition queries. However, as the number of conditions increases, their performance declines more sharply. As shown in Table 1, for Query1, rerankers almost outperformed all retrieval models. bge-reranker-v2-gemma achieves a Win Rate of 91.07% at Query1 but drops to 56.09% at Query10, a drastic 34.98% decrease. Similarly, bge-reranker-v2-m3 declines from 87.14% to 44.87%, a drop of 42.27%, and followIR shows a drop from 83.41% to 43.52%, with a drop of 39.89%. In contrast, retrievers exhibit an average Win Rate decline of 14.06%.
Sparse retrieval models show performance improvement
Unlike dense retrievers and rerankers, BM25’s performance improves with query complexity, increasing from 28.59% (Query1) to 39.87% (Query10). This is likely due to its lexical matching mechanism, where more query conditions lead to greater lexical overlap with candidate documents, enhancing retrieval effectiveness. However, BM25’s overall win rate remains lower than dense retrievers, and positives rank below HNs (win rate < 0.5), suggesting that hard negatives consistently outrank positives. This may stem from dataset construction, where queries and HNs were generated or paraphrased by GPT-4o, potentially introducing lexical bias [^4] Appendix E further explores challenges in using generated datasets..
4.2 Results for Relevance Monotonicity
Fig. 4 illustrates the trend of average Win Rate in the multi-condition retrieval setting of Task 2555The complete results are provided in Appendix D.2, which evaluates the model’s ability to distinguish the relevance hierarchy among documents with varying conditions. Several key observations can be made:
Relevance monotonicity failure
As documents become increasingly hard (i.e., satisfying more conditions in the query), retrieval and reranking models struggle to distinguish between $d_{k}$ and $d_{k-1}$, leading to a decline in Win Rate performance. This failure emphasizes the challenge of preserving relevance monotonicity in multi-condition retrieval settings and highlights a gap in current model capabilities when handling complex queries.
Sensitive to “exact match” and “complete mismatches”
We observe a slight upward trend at the end of Win Rate curves for most dense retrievers, likely due to their contrastive learning-based training. Traditional contrastive learning treats retrieval as a binary task, pulling query-positive pairs closer while pushing negatives further apart, without accounting for partial matches. As a result, dense retrievers perform more reliably in clear-cut “exact match” or “complete mismatch” cases.
4.3 Results for Format Invariance
Model People Books Movies Medical Legal Avg.
Spare Retriever
BM25 14.86 17.88 16.84 19.25 12.14 16.19
Dense Retriever
jina-embeddings-v3 10.55 8.65 10.24 14.72 13.10 11.45
gte-large-en-v1.5 11.74 8.84 12.96 15.70 15.16 12.88
NV-Embed-v2 10.17 8.80 7.52 10.17 8.94 9.12
bge-en-icl 13.48 12.74 15.18 19.81 14.44 15.13
gte-Qwen2-7B-instruct 12.71 15.56 13.62 16.37 17.56 15.16
gte-Qwen2-1.5B-instruct 12.38 13.26 10.48 16.81 12.51 13.09
e5-mistral-7b-instruct 9.17 9.92 8.20 10.75 12.25 10.06
GritLM-7B 8.52 5.35 8.32 8.98 9.86 8.21
LLM2Vec 12.81 7.93 9.56 8.12 10.49 9.78
Reranker
bge-reranker-v2-m3 42.40 32.22 34.82 28.35 31.24 33.81
bge-reranker-v2-gemma 27.50 18.94 16.52 13.42 24.41 20.16
followIR 35.81 31.60 23.70 25.07 28.43 28.92
Table 2: Flip Rate for query format shift (Task 3). The Flip Rate reflects the win rate reversal when switching the query format from instruction-style to descriptive-style.
Table 2 presents the Flip Rate induced by query format variations. Most models exceed 10%, indicating a substantial impact of query formulation on retrieval performance. BM25 records a 16.19% Flip Rate, which can serve as a baseline for the lexical differences caused by query format variations.
Dense retrieval models show relatively lower sensitivity, with Flip Rates between 8% to 16%. GritLM-7B (8.21%), NV-Embed-v2 (9.12%), and LLM2Vec (9.78%) exhibit less variation, suggesting higher robustness to query styles.
In contrast, reranking models show significantly higher sensitivity to query format changes, with Flip Rates exceeding 20%. The highest Flip Rate observed is 33.81% for bge-reranker-v2-m3, confirming their susceptibility to query formulation changes.
(a) Retrieval Performance when padding to 512 words
(b) Retrieval Performance when padding to 512 words
Figure 5: Retrieval performance when padding the document set to 512 words and 1024 words.
5 Analysis
5.1 Retrievers vs. Rankers
Our experiments reveal notable differences between retrieval and reranking models across the three tasks. Specifically, retrieval models demonstrate greater robustness to query complexity, better preservation of relevance monotonicity, and stronger invariance to query format changes compared to rerankers. We analyze these differences through the lens of relevance computation and attention mechanisms.
Retrieval models: robustness induced by bidirectional attention and dual-encoder.
Retrieval models typically employ a dual-encoder architecture, where queries and documents are independently encoded before computing their similarity using dot-product or cosine similarity. This independent computation ensures that the generation of query and document embeddings remains unaffected by each other. At the same time, bidirectional attention enables the model to better capture the overall semantic meaning of the query.
Another critical factor is the training datasets. Many dense retrieval models are trained on a mixture of retrieval-specific and general textual datasets (Lee et al., 2024; BehnamGhader et al., 2024; Wang et al., 2024a), which enhances their generalization across diverse retrieval scenarios and query styles. This, in turn, improves their robustness against query complexity.
Reranking models: sensitivity to query complexity.
Figure 6: Attention heatmap of cross-encoder model (bge-reranker-m3): effect of increasing query conditions. When query conditions are few (1-3), the cross-encoder effectively attends to relevant conditions in the document, as the number of conditions increases, attention becomes more diluted.
Reranking models compute relevance by jointly processing the query and document, primarily through: (1) cross-encoders, which perform token-level relevance comparison through cross-attention (e.g., bge-reranker-v2-m3), and (2) generative models that estimate relevance using LLM-based agents (e.g., bge-reranker-v2-gemma and FollowIR). Both architectures rely on deep query-document interaction, making them more sensitive to input complexity, such as changes in condition quantity and query format.
Our experiments show that rerankers outperform retrieval models in ranking effectiveness for single-condition queries, suggesting their advantage in capturing fine-grained query-document relevance for short and simple queries. However, as query complexity increases, their performance declines more sharply, eventually falling behind retrieval models. Table 1 shows that rerankers’ average Win Rate drops by 39.05%, more than twice the decline of retrieval models (14.06%).
A case study in Fig. 6 provides further insight into this decline by analyzing the attention heatmap of a cross-encoder reranker. When the query contains only a few conditions (e.g., 1–3), the model effectively aligns each condition with relevant document sentences. However, as conditions increase, attention becomes unevenly distributed, with earlier conditions receiving more focus while later ones are overlooked. This imbalance likely contributes to the sharp drop in Win Rate observed in Task 2—as later conditions receive insufficient attention, the model struggles to distinguish relevance variations in corresponding document sentences.
These findings highlight a key limitation in existing reranking models: their reliance on fine-grained query-document interactions becomes a bottleneck in multi-condition scenarios. While reranking models outperform retrieval models for queries with fewer conditions, their effectiveness diminishes as query complexity grows.
5.2 The Effect of Document Side
Previous experiments focused on query-side variations. We now investigate how changes on the document side, particularly document length, affect retrieval effectiveness.
Figure 5 shows the impact of document length (padded to 512 and 1024 words) on relevance monotonicity [^6] The complete results and the padding filler are provided in Appendix D.3. Compared to original documents (Fig.4), increasing document length does not change the overall trend in Task 2—as hard negatives (HNs) become more challenging, models struggle more to distinguish relevance between documents satisfying $k$ vs. $k-1$ conditions. This decline is more pronounced when document length increases to 1024 words.
Rerankers are particularly affected, showing a steeper win rate drop than retrieval models. Since they process queries and documents jointly, longer inputs likely increase attention dispersion in cross-encoders and complexity in generative models, weakening relevance estimation.
Among retrieval models, GritLM and Nv-embed demonstrate strong robustness to document length variations. At 512 words, GritLM (66.66%) and Nv-embed (62.91%) achieve the highest average win rate. When document length is increased to 1024 words, Nv-embed (63.41%) outperforms GritLM (59.46%), indicating its ability to retain essential information in longer documents. This suggests that Nv-embed is more effective at capturing key content even when document length increases.
5.3 Robustness of GritLM and NV-Embed
(a) Mean pooling
(b) <EOS> pooling
(c) Latent attention layer pooling
Figure 7: Impact of condition position on different pooling methods. The condition is placed at different positions (1–10) in the query, with other positions filled by ten “[unused0]” tokens (example from Table 9). The dashed line represents the original data, while the solid line shows the Gaussian-smoothed trend (kernel size = 1) for clarity.
GritLM and hybrid attention.
GritLM exhibits strong robustness in multi-condition retrieval, which we hypothesize is due to its hybrid attention mechanism, integrating causal attention for generative task and bidirectional attention for embedding learning.
Recent work has incorporated bidirectional attention into LLM-based retrievers, such as LLM2Vec (BehnamGhader et al., 2024) and Nv-Embed (Lee et al., 2024), which discard causal attention after converting LLMs into embedding models. In contrast, GritLM preserves causal attention, maintaining generative reasoning capabilities alongside representation learning. This likely explains its superior adaptability in multi-condition retrieval. Empirical results further validate this advantage: GritLM achieves state-of-the-art performance in complex instruction-following retrieval (Zhou et al., 2025) and ranks second in reasoning-intensive retrieval tasks (Su et al., 2024), reinforcing the importance of hybrid attention in embedding-based retrieval.
Nv-Embed and latent attention layer pooling.
Nv-Embed performs suboptimally in most multi-condition retrieval tasks, which we hypothesize is due to its Latent Attention Layer Pooling strategy. Common embedding methods include: 1) mean pooling, and 2) the last <EOS> token embedding. Mean pooling averages all token representations to generate a global representation, but potentially dilute critical information. Last <EOS> token embedding relies solely on the final token, which could introduce recency bias.
A case study in Fig. 7 examines how different pooling methods handle condition placement in queries. The results show that mean pooling leads to decreasing similarity scores as the condition moves further back, while <EOS> pooling exhibits an increasing trend, confirming its recency bias. In contrast, latent attention layer pooling maintains relatively stable performance across different condition positions, indicating improved information retention.
We validated the robustness of Nv-Embed in handling longer documents in Section 5.2. Notably, at 1024 words, it maintains consistent monotonic relevance, preserving key information more effectively than other models. These findings suggest that its latent attention layer pooling potentially mitigates information loss, enhancing retrieval stability in multi-condition scenarios.
6 Conclusion
In this paper, we introduce MultiConIR, the first benchmark specifically designed to evaluate retrieval models in multi-condition scenarios. Unlike existing datasets that primarily focus on single-condition queries from search engines, MultiConIR captures real-world complexity by incorporating five diverse domains: books, movies, people, medical cases, and legal documents.
We propose three tasks to systematically assess retrieval and reranking models on multi-condition robustness, monotonic relevance ranking, and query format sensitivity. Through extensive experiments, we analyze model performance across architectures and sizes, examining their ability to distinguish fine-grained conditions, preserve semantic information, and adapt to contextual variations.
Our findings reveal that existing retrieval and reranking models struggle with multi-condition retrieval, with rerankers suffering severe performance degradation as query complexity increases. We further investigate the performance gap between retrieval and reranking models, exploring potential reasons for these discrepancies. Finally, we highlight the strengths of GritLM and Nv-Embed, which demonstrate enhanced adaptability to multi-condition queries, offering insights for future retrieval models.
Limitations
While MultiConIR provides a novel benchmark for evaluating retrieval models in multi-condition scenarios, several limitations should be acknowledged. First, our dataset relies on automated query generation and hard negative creation, which may introduce biases in condition representation despite efforts to ensure accuracy. These biases could affect retrieval models’ ability to distinguish fine-grained differences. Second, our evaluation focuses on retrieval tasks and does not cover reasoning-based retrieval or interactive search scenarios. Real-world systems often incorporate reranking, user feedback, and hybrid retrieval, which are not explicitly modeled. Lastly, our dataset does not fully consider query reformulation strategies or multi-turn retrieval, limiting its applicability to dynamic search environments. These limitations highlight the need for further research into multi-condition retrieval, particularly in addressing dataset biases, expanding evaluation scopes, and integrating retrieval with realistic user interactions.
Ethics Statement
This study adheres to ethical standards in AI research, ensuring transparency and reproducibility in dataset construction and model evaluation while exclusively using publicly available pre-trained models for experiments. Dataset Considerations: MultiConIR is built from publicly available sources and does not contain sensitive or personally identifiable information. Given its inclusion of medical and legal documents, we apply strict data filtering and safety measures to respect model safety constraints and prevent the generation of harmful or misleading content. Additionally, we recognize that automatically generated queries and hard negatives may introduce biases. Therefore, during dataset construction, we take measures to minimize the impact of inherent language model biases on retrieval tasks. MultiConIR aims to advance multi-condition retrieval research while ensuring data fairness and ethical compliance. We encourage future research to further explore bias detection strategies in retrieval dataset, enhancing model fairness and reliability in diverse corpus environments.
Appendix A Details of Models
For each model used in this paper, Table 3 provides details on model size, architecture, maximum input context length, and whether instructions is included.
Model Size Architecture Instruction Max length Pooling Method
Sparse Retriever
BM25 (Robertson and Zaragoza, 2009) N/A Sparse No N/A N/A
Dense Retriever
jina-embeddings-v3 (Sturua et al., 2024) 572M Encoder No 4K Mean
gte-large-en-v1.5 (Li et al., 2023b) 434M Encoder No 8K <GLS>
NV-Embed-v2 (Lee et al., 2024) 7.8B Decoder Yes 32K Latent Attention Layer
bge-en-icl (Li et al., 2024) 7.1B Decoder Yes 32K <EOS>
gte-Qwen2-7B-instruct (Li et al., 2023b) 7.6B Decoder Yes 131K <EOS>
gte-Qwen2-1.5B-instruct (Li et al., 2023b) 1.5B Decoder Yes 131K <EOS>
e5-mistral-7b-instruct (Wang et al., 2024a) 7.1B Decoder Yes 32K <EOS>
GritLM-7B (Muennighoff et al., 2024) 7.2B Decoder Yes 4K Mean
LLM2Vec (BehnamGhader et al., 2024) 7.5B Decoder Yes 8K Mean
Reranker
bge-reranker-v2-m3 (Chen et al., 2024) 568M Cross-Encoder No 8k N/A
bge-reranker-v2-gemma (Chen et al., 2024) 2.5B Decoder Yes 1k N/A
followIR (Weller et al., 2024a) 7.2B Decoder Yes 4k N/A
Table 3: Details of models used in experiments. We list the number of parameters of each model except the sparse model (BM25). Regarding the model architecture, we distinguish between sparse models, dense models, and rerankers. Dense models are further classified as Encoders or Decoders. Rerankers are categorized into Cross Encoders and Decoders (LLM-based generative relevance scoring). The Instruction column indicates whether instructions are included in the retrieval process. Max length denotes the maximum input length used for each model in the experiments. The Pooling Method represents the approach used by the model to obtain embeddings.
For Dense Retrieval models that require instructions (NV-Embed-v2, bge-en-icl, gte-Qwen2-7B-instruct, gte-Qwen2-1.5B-instruct, e5-mistral-7b-instruct, GritLM-7B, and LLM2Vec), we use the following instruction:
“Given a domain retrieval query, retrieve documents that meet the specified conditions.”
For LLM-based rerankers (bge-reranker-v2-gemma and followIR), we adopt the model’s default prompt. For example, bge-reranker-v2-gemma uses the following prompt:
“Given a query A and a passage B, determine whether the passage contains an answer to the query by providing a prediction of either ‘Yes’ or ‘No’.”
For models that do not require instructions, we directly input the query and document, such as jina-embeddings-v3, gte-large-en-v1.5, and bge-reranker-v2-m3.
Appendix B Prompt Templates For Constructing MultiCon Dataset
Table 4, 5, and 6 present the prompts used in Steps 1 to 3 for constructing our MultiCon dataset.
[…]
Task Prompt
Step 1: Condition Sentence Extraction I will provide you a document of {domain}, you should extract ten detailed sentences that represent the key conditions the document satisfies.
Please adhere to the following guidelines:
- Extract fine-grained condition-related sentences relevant to {domain}, such as {domain_features}.
- Do not paraphrase; use the original sentences from the document.
- Ensure each sentence is semantically intact and not conflict with the context.
- Format the output as an array, e.g., [“sentence1”, “sentence2”, …, “sentence10”].
Here is the document: {domain_document}.
Return array only.
Table 4: Prompt for GPT-4o to extract condition sentence (Step 1).
Task Prompt
Step 2: Query Generation (Instruction-style) I will provide you {num} condition-related sentences; formulate a retrieval query for me.
Here are a few examples for reference:
- {Instruction-style example 1}
- {Instruction-style example 2}
Please adhere to the following guidelines:
- Each sentence represents a condition; with {num} sentences, the number of conditions is {num}.
- The query should be instruction-style, explicitly listing all conditions. - Each condition should be around 10 words.
- Make conditions concise, summarizing each sentence.
- You can paraphrase and modify keywords while maintaining meaning.
Here are the sentences: {info}.
Return one query only. Do not include extra information.
Step 2: Query Generation (Descriptive-style) I will provide you {num} condition-related sentences; formulate a retrieval query for me.
Here are a few examples for reference:
- {Descriptive-style example 1}
- {Descriptive-style example 2}
Please adhere to the following guidelines:
- Each sentence represents a condition; with {num} sentences, the number of conditions is {num}.
- The query should be descriptive-style, integrating and describing all conditions in natural language.
- Each condition should be around 10 words.
- Make conditions concise, summarizing each sentence.
- You can paraphrase and modify keywords while maintaining meaning.
Here are the sentences: {info}.
Return one query only. Do not include extra information.
Table 5: Prompt for GPT-4o to generate queries with varying conditions (Step 2).
Task Prompt
Step 3: Hard Negative Sentence Making (For Books, Movies, Medical Case, and Legal Document Datasets) I will provide you one query and one sentence, generate a modified sentence for me.
Here are a few examples for reference:
Query: - {query}
Sentence: - {condition sentence}
Modified: - {hard negative sentence}
Please adhere to the following guidelines:
- Modify the sentence so that its meaning no longer aligns with the query.
- Keep key terms unchanged.
- Ensure the new sentence is semantically different from the original.
Here is the query: {query}.
Here is the Sentence: {information}.
Return only the modified sentence.
Step 3: Hard Negative Sentence Making (For People Dataset) I will provide you one query and one sentence, generate a modified sentence for me.
Here are a few examples for reference:
Query: - Who is the American artist that went to RISD?
Sentence: - He went to RISD for graduate school.
Modified: - He went to ACCA for graduate school, but his sister went to RISD.
Please adhere to the following guidelines:
- Modify the sentence so that its meaning no longer aligns with the query.
- Keep key terms unchanged, but introduce dummy information to mislead the retrieval model. For example, if the original sentence states, “He went to RISD for graduate school,” you can modify it to, “He went to ACCA for graduate school, but his sister went to RISD,” where the key term (RISD) remains but is assigned to an irrelevant entity (his sister).
- Ensure the new sentence is semantically different from the original by using different wording and synonymous substitution.
- The changed sentence should prevent the query from retrieving it as relevant information.
Here is the query: {query}.
Here is the sentence: {information}.
Return only the modified sentence.
Table 6: Prompt for GPT-4o to modify the condition sentence to hard negative sentence (Step 3).
Appendix C Details And Examples Of The MultiCon Dataset
Table 7 presents the five domains and their source data used for constructing our MultiCon dataset. Tables 9, 10, 8, 11, and 12 illustrate examples from their respective domains.
Domain Records Source Data
People 420 People Wikipedia Dat(Mahajan, 2017)
Books 482 Books Dataset (Rustamov, 2021)
Movies 500 Wikipedia Movie Plots (Robischon, 2018)
Medical Case 479 Medical Cases (HPE AI Solutions, 2023)
Legal Document 426 LexGLUE (Chalkidis et al., 2022)
Table 7: Domain-specific Source Data References. The Records column indicates the number of data entries for each domain in the MultiCon dataset.
Query 10 Positive HN 1
Find a notable individual who meets these criteria: 1. Studies plasma melatonin to assess biological rhythm disorders. 2. Identified 25-hour circadian rhythms in totally blind individuals. 3. Worked at NIMH in Bethesda, Maryland before 1981. 4. Collaborated closely with Robert Sack in Oregon. 5. Researches chronobiologic sleep and mood disorders. 6. Graduated from University of Chicago in nineteen seventies with MD and PhD. 7. Had around 90 publications on PubMed by 2005. 8. Is a full professor and vice-chair of Psychiatry at OHSU. 9. Focuses on bright light exposure and melatonin treatments. 10. Studies disorders like winter depression, jet lag, and shift work maladaptation. Relying on a very precise assay for plasma melatonin, a hormone that has a clearly defined 24-hour pattern of secretion, biological rhythm disorders can be assessed and their treatment can be monitored. Totally blind individuals have 25-hour circadian rhythms, drifting an hour later each day unless they take a melatonin capsule at a certain time every day. Prior to moving to Oregon in 1981, Lewy was at the National Institute of Mental Health (NIMH) in Bethesda, Maryland, working with senior colleague Thomas Wehr. In Oregon, he has worked closely with Robert L. Sack. He describes his research as follows: ’My laboratory studies chronobiologic sleep and mood disorders.’ Alfred J. Lewy, aka Sandy Lewy, graduated from University of Chicago in 1973 after studying psychiatry, pharmacology, and ophthalmology. As of December 2005, he had 94 publications available on PubMed. He is a full professor and vice-chair of the Department of Psychiatry at OHSU (Oregon Health Science University) and holds an MD and PhD. Current research is focused on developing bright light exposure and melatonin administration as treatment modalities for these disorders. These disorders include winter depression, jet lag, maladaptation to shift work, and certain types of sleep disturbances. Relying on a very precise assay for plasma melatonin, a hormone that has a clearly defined 24-hour pattern of secretion, biological rhythm disorders can be assessed and their treatment can be monitored. Totally blind individuals have 25-hour circadian rhythms, drifting an hour later each day unless they take a melatonin capsule at a certain time every day. Prior to moving to Oregon in 1981, Lewy was at the National Institute of Mental Health (NIMH) in Bethesda, Maryland, working with senior colleague Thomas Wehr. In Oregon, he has worked closely with Robert L. Sack. He describes his research as follows: ’My laboratory studies chronobiologic sleep and mood disorders.’ Alfred J. Lewy, aka Sandy Lewy, graduated from University of Chicago in 1973 after studying psychiatry, pharmacology, and ophthalmology. As of December 2005, he had 94 publications available on PubMed. He is a full professor and vice-chair of the Department of Psychiatry at OHSU (Oregon Health Science University) and holds an MD and PhD. Current research is focused on developing bright light exposure and melatonin administration as treatment modalities for these disorders. These lifestyle changes address issues like winter depression, jet lag, shift work maladaptation, and certain types of sleep disturbances.
Table 8: An example in domain of People
Query 10 Positive HN 1
Find a notable individual who meets these criteria: 1. Details Bernie Madoff’s \$65B Ponzi collapse. 2. Covers the impact on national media. 3. Investigates Madoff’s history of fraud. 4. Offers deep insight into Madoff’s family. 5. Contains exclusive news and material. 6. Includes Bernie’s Little Black Book. 7. Features Madoff’s past calendars. 8. Provides details on the origins of his scam. 9. Includes cooperation from Eleanor Squillari. 10. Reveals SEC connections to Madoff. The collapse of Bernie Madoff’s Ponzi scheme led to the instant evaporation of \$65 billion of wealth. The effects of Madoff’s brazen fraud were felt most closely in New York and Palm Beach but the story was, and continues to be, front page news across the country. Brian Ross and his team of investigators shed an unyielding light onto Madoff’s scheme–how he got started, how he succeed for so long, who helped him, and who shielded him from early investigations. This is an incisive and voyeuristic look into this first family of financial crime. The Madoff Chronicles includes a vast array of news and material that readers won’t find anywhere else. Contains a reproduction of Bernie’s Little Black Book. Ross has also secured Madoff’s calendar for the past three years and other never-before-seen documents from inside the Madoff empire, straight from his desk. Read key details of how Madoff carried out his scam and the revelation that he began the fraud from almost the first day, in the 1960s. Extensive cooperation by Madoff’s personal assistant, Eleanor Squillari. Contains incriminating connections between Madoff and certain members of the SEC. The collapse of Bernie Madoff’s Ponzi scheme led to the instant evaporation of \$65 billion of wealth. The effects of Madoff’s brazen fraud were felt most closely in New York and Palm Beach but the story was, and continues to be, front page news across the country. Brian Ross and his team of investigators shed an unyielding light onto Madoff’s scheme–how he got started, how he succeed for so long, who helped him, and who shielded him from early investigations. This is an incisive and voyeuristic look into this first family of financial crime. The Madoff Chronicles includes a vast array of news and material that readers won’t find anywhere else. Contains a reproduction of Bernie’s Little Black Book. Ross has also secured Madoff’s calendar for the past three years and other never-before-seen documents from inside the Madoff empire, straight from his desk. Read key details of how Madoff carried out his scam and the revelation that he began the fraud from almost the first day, in the 1960s. Extensive cooperation by Madoff’s personal assistant, Eleanor Squillari. Contains no documented communications between Madoff and SEC members.
Table 9: An example in domain of Book
Query 10 Positive HN 1
Find a movie that matches all conditions: 1. Originated from American. 2. Plot: Mrs. Lowe and Black were lovers. 3. Plot: Terry carelessly spends money. 4. Plot: Terry promoted to ship foreman. 5. Cast: Mary Miles Minter, Allan Forrest. 6. Plot: Julia Deep works behind exchange desk. 7. Director: Lloyd Ingraham. 8. Plot: Terry’s spending takes a toll. 9. Plot: Lottie sees Terry and Julia at park. 10. Release Year: 1918. Origin/Ethnicity: American
Meanwhile, it is revealed Mrs. Lowe and Black were once lovers. He is spending his money carelessly and doesn’t put any time in paying the bills, much to the dislike of the department store owner Timothy Black. Soon, Terry is promoted to a foreman on a ship.
Cast: Mary Miles Minter, Allan Forrest Julia Deep is a young woman working behind the exchange desk at a department store.
Director: Lloyd Ingraham
After a while, Terry’s money spending takes its toll. Lottie gets distracted and does not notice Terry and Julia at the park.
Release Year: 1918
Origin/Ethnicity: American
Meanwhile, it is revealed Mrs. Lowe and Black were once lovers. He is spending his money carelessly and doesn’t put any time in paying the bills, much to the dislike of the department store owner Timothy Black. Soon, Terry is promoted to a foreman on a ship.
Cast: Mary Miles Minter, Allan Forrest Julia Deep is a young woman working behind the exchange desk at a department store.
Director: Lloyd Ingraham
Eventually, Terry’s frugality leads to financial growth. Lottie gets distracted and does not notice Terry and Julia at the park.
Release Year: 1918
Table 10: An example in domain of Movie
Query 10 Positive HN 1
Find a case where the patient: 1. Underwent ascending aortic arch angiogram. 2. Had left common carotid artery angiogram. 3. Received right common carotid artery angiogram. 4. Undergone left subclavian artery angiogram. 5. Had right iliac angiogram with runoff. 6. Performed bilateral cerebral angiograms. 7. Experienced TIA and moderate carotid stenosis. 8. Had 400 ml blood loss. 9. Provided informed consent for the procedure. 10. Received a 6-French sheath in the right femoral artery. PROCEDURE PERFORMED:
1. Selective ascending aortic arch angiogram. 2. Selective left common carotid artery angiogram. 3. Selective right common carotid artery angiogram. 4. Selective left subclavian artery angiogram. 5. Right iliac angio with runoff. 6. Bilateral cerebral angiograms were performed as well via right and left common carotid artery injections.
INDICATIONS FOR PROCEDURE: TIA, aortic stenosis, postoperative procedure. Moderate carotid artery stenosis.
ESTIMATED BLOOD LOSS: 400 ml.
After obtaining informed consent, the patient was brought to the cardiac catheterization suite in postabsorptive and nonsedated state. Using modified Seldinger technique, a 6-French sheath was placed into the right common femoral artery and vein without complication.
PROCEDURE PERFORMED:
1. Selective ascending aortic arch angiogram. 2. Selective left common carotid artery angiogram. 3. Selective right common carotid artery angiogram. 4. Selective left subclavian artery angiogram. 5. Right iliac angio with runoff. 6. Bilateral cerebral angiograms were performed as well via right and left common carotid artery injections.
INDICATIONS FOR PROCEDURE: TIA, aortic stenosis, postoperative procedure. Moderate carotid artery stenosis.
ESTIMATED BLOOD LOSS: 400 ml.
After obtaining informed consent, the patient was brought to the cardiac catheterization suite in postabsorptive and nonsedated state. A 6-French sheath was used in the left femoral artery and vein with minor complications, employing the modified Seldinger technique.
Table 11: An example in domain of Medical Case
Query 10 Positive HN 1
Find a case where:
1. Michigan Legislature enacted a statute in 1987.
2. Petitioners challenged the statute under Contract Clause and Due Process Clause.
3. The statute affected workers injured before March 31, 1982.
4. Petitioners argued a 1981 law allowed reduction of workers’ compensation benefits.
5. The Michigan Supreme Court accepted petitioners’ interpretation in 1985.
6. Legislature introduced a bill to overturn the court’s decision.
7. House Bill 5084 was introduced in October 1985.
8. The bill became law on May 14, 1987.
9. Petitioners were ordered to refund nearly \$25 million.
10. Michigan Supreme Court upheld the statute for lacking vested rights and rational purpose.
In 1987, the Michigan Legislature enacted a statute that had the effect of requiring petitioners General Motors Corporation (GM) and Ford Motor Company (Ford) to repay workers’ compensation benefits GM and Ford had withheld in reliance on a 1981 workers’ compensation statute. Petitioners challenge the provision of the statute mandating these retroactive payments on the ground that it violates the Contract Clause and the Due Process Clause of the Federal Constitution. The benefit coordination provision did not specify whether it was to be applied to workers injured before its effective date, March 31, 1982. Petitioners took the position that the 1981 law allowed them to reduce workers’ compensation benefits to workers injured before March 31, 1982, who were receiving benefits from other sources. In 1985, petitioners’ interpretation was accepted by the Michigan Supreme Court. Chambers v. General Motors Corp., decided together with Franks v. White Pine Copper Div., Copper Range Co., 422 Mich. 636, 375 N.W.2d 715. The Michigan Legislature responded almost immediately by introducing legislation to overturn the court’s decision. On October 16, 1985, before the Michigan Supreme Court had ruled on the motion for rehearing in Chambers, House Bill 5084 was introduced. The amended Senate bill passed into law on May 14, 1987. 1987 Mich.Pub.Acts No. 28. As a result of the 1987 statute, petitioners were ordered to refund nearly \$25 million to disabled employees. The Michigan Supreme Court upheld the statute against these challenges, on the ground that the employers had no vested rights in coordination for Contract Clause purposes, and that the retroactive provisions furthered a rational legislative purpose. 436 Mich. 515, 462 N.W.2d 555 (1990). In 1987, the Michigan Legislature enacted a statute that had the effect of requiring petitioners General Motors Corporation (GM) and Ford Motor Company (Ford) to repay workers’ compensation benefits GM and Ford had withheld in reliance on a 1981 workers’ compensation statute. Petitioners challenge the provision of the statute mandating these retroactive payments on the ground that it violates the Contract Clause and the Due Process Clause of the Federal Constitution. The benefit coordination provision did not specify whether it was to be applied to workers injured before its effective date, March 31, 1982. Petitioners took the position that the 1981 law allowed them to reduce workers’ compensation benefits to workers injured before March 31, 1982, who were receiving benefits from other sources. In 1985, petitioners’ interpretation was accepted by the Michigan Supreme Court. Chambers v. General Motors Corp., decided together with Franks v. White Pine Copper Div., Copper Range Co., 422 Mich. 636, 375 N.W.2d 715. The Michigan Legislature responded almost immediately by introducing legislation to overturn the court’s decision. On October 16, 1985, before the Michigan Supreme Court had ruled on the motion for rehearing in Chambers, House Bill 5084 was introduced. The amended Senate bill passed into law on May 14, 1987. 1987 Mich.Pub.Acts No. 28. As a result of the 1987 statute, petitioners were ordered to refund nearly \$25 million to disabled employees. The Michigan Supreme Court found the statute invalid on the grounds that the retroactive provisions did not further a rational legislative purpose and that the employers had vested rights in coordination for Contract Clause purposes. 436 Mich. 515, 462 N.W.2d 555 (1990).
Table 12: An example in domain of Legal Document
Appendix D Complete Results
D.1 Complete Results Of Task1
Table 13, 14, 15, 16, and 17 present the complete results of Task 1 for the five datasets: People, Books, Movies, Medical Case, and Legal Document.
Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 Decline
Spare Retriever
BM25 23.10 25.24 22.62 26.43 26.19 23.57 24.05 22.14 22.38 23.81 -0.71
Dense Retriever
jina-embeddings-v3 88.81 78.33 78.33 76.90 72.14 68.57 67.38 66.19 67.62 65.95 22.86
gte-large-en-v1.5 72.14 76.19 73.10 68.57 69.29 61.43 66.19 62.38 69.05 63.81 8.33
NV-Embed-v2 82.14 89.52 85.48 85.48 87.38 84.05 87.38 84.05 84.05 86.67 -4.52
bge-en-icl 85.00 85.95 86.43 85.00 83.81 84.29 86.43 87.14 85.48 83.81 1.19
gte-Qwen2-7B-instruct 58.10 70.95 69.05 70.71 71.90 71.43 71.43 69.76 72.62 69.76 -11.67
gte-Qwen2-1.5B-instruct 61.90 74.76 70.24 71.19 73.57 70.48 73.81 66.90 65.71 65.00 -3.10
e5-mistral-7b-instruct 76.43 77.86 73.33 75.48 68.10 71.43 67.62 68.81 69.05 73.57 2.86
GritLM-7B 79.05 84.52 80.71 81.67 81.67 77.86 83.10 82.62 81.43 82.62 -3.57
LLM2Vec 79.05 74.76 71.90 75.71 74.05 75.95 75.00 78.81 76.19 80.00 -0.95
Reranker
bge-reranker-v2-m3 90.00 85.95 78.57 73.81 74.29 62.14 63.57 56.67 52.14 36.67 53.33
bge-reranker-v2-gemma 92.62 93.33 91.67 88.33 89.29 84.76 81.43 72.38 69.52 50.71 41.90
followIR 84.52 80.48 76.43 68.57 60.71 56.90 50.00 34.05 20.95 8.57 75.95
Table 13: Win Rate Score as the Number of Conditions in the Query Increases on the People Dataset (Task 1)
Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 decline
Spare Retriever
BM25 45.85 47.93 47.93 49.17 47.72 51.24 50.21 48.34 53.94 46.06 -0.21
Dense Retriever
jina-embeddings-v3 76.14 72.41 71.99 71.37 67.01 64.73 67.43 63.07 63.28 58.51 17.63
gte-large-en-v1.5 73.86 75.93 71.99 70.33 68.67 69.71 65.77 66.60 64.32 70.95 2.90
NV-Embed-v2 78.01 73.65 75.93 68.88 69.71 69.50 68.46 68.05 67.84 71.58 6.43
bge-en-icl 78.42 76.76 75.31 72.20 72.41 69.29 70.54 70.12 70.95 72.41 6.02
gte-Qwen2-7B-instruct 71.99 63.90 64.52 61.83 54.77 50.00 50.00 49.17 48.55 50.83 21.16
gte-Qwen2-1.5B-instruct 70.54 69.09 70.33 69.92 65.98 60.58 64.11 62.45 64.11 56.85 13.69
e5-mistral-7b-instruct 73.03 65.77 68.88 65.98 60.58 59.96 59.34 60.58 60.37 68.67 4.36
GritLM-7B 80.91 74.69 81.33 73.86 75.73 73.86 79.25 77.80 78.01 80.50 0.41
LLM2Vec 80.08 70.12 72.41 70.12 68.05 66.80 64.32 65.35 65.15 64.52 15.56
Reranker
bge-reranker-v2-m3 83.20 79.05 76.35 71.99 72.20 72.82 74.48 68.67 60.58 51.66 31.54
bge-reranker-v2-gemma 87.76 84.44 82.78 81.74 80.08 75.31 74.48 72.82 70.75 70.54 17.22
followIR 85.89 78.22 77.59 72.41 68.67 68.46 61.41 60.37 53.11 47.30 38.59
Table 14: Win Rate Score as the Number of Conditions in the Query Increases on the Books Dataset (Task 1)
Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 decline
Spare Retriever
BM25 28.60 34.60 33.60 37.20 34.80 36.80 37.00 38.40 37.60 39.00 -10.40
Dense Retriever
jina-embeddings-v3 77.60 77.00 75.20 75.40 70.80 72.00 69.40 68.40 71.80 67.20 10.40
gte-large-en-v1.5 78.40 81.00 76.60 75.60 73.80 71.00 71.00 69.00 67.20 59.40 19.00
NV-Embed-v2 81.00 81.80 80.40 78.60 80.00 76.00 77.40 76.60 72.20 67.00 14.00
bge-en-icl 84.40 83.20 80.60 79.20 78.80 77.00 76.00 75.80 70.40 66.60 17.80
gte-Qwen2-7B-instruct 74.60 80.00 74.60 72.80 72.40 72.00 68.80 75.00 67.60 67.40 7.20
gte-Qwen2-1.5B-instruct 80.40 83.00 76.00 74.80 73.00 73.80 71.00 74.60 72.20 68.40 12.00
e5-mistral-7b-instruct 73.40 71.80 66.00 64.80 70.00 61.20 65.60 62.80 63.00 55.60 17.80
GritLM-7B 85.00 83.60 80.00 80.80 79.80 72.80 76.40 72.60 71.40 73.80 11.20
LLM2Vec 87.20 89.40 86.40 84.40 81.80 81.60 79.60 77.80 75.40 74.20 13.00
Reranker
bge-reranker-v2-m3 92.80 90.80 80.20 81.60 79.20 75.60 71.00 63.40 67.00 60.40 32.40
bge-reranker-v2-gemma 92.20 91.60 86.00 89.20 83.40 79.40 77.60 78.20 70.80 61.60 30.60
followIR 92.60 86.40 83.40 88.00 81.00 81.00 83.40 81.40 82.60 85.00 7.60
Table 15: Win Rate Score as the Number of Conditions in the Query Increases on the Movies Dataset (Task 1)
Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 Decline
Spare Retriever
BM25 25.68 28.18 31.52 32.99 37.37 36.33 35.49 39.25 37.16 40.71 -15.03
Dense Retriever
jina-embeddings-v3 68.89 64.93 65.34 65.55 55.32 59.08 56.16 63.67 60.54 59.29 9.60
gte-large-en-v1.5 77.24 81.84 76.83 75.57 71.82 70.56 70.35 67.01 68.48 62.42 14.82
NV-Embed-v2 77.24 73.07 74.53 73.70 65.97 62.63 67.43 68.68 67.22 63.67 13.57
bge-en-icl 80.79 78.91 76.41 77.04 68.89 68.68 72.23 71.19 69.31 67.64 13.15
gte-Qwen2-7B-instruct 71.61 73.70 73.49 72.65 66.60 67.64 66.18 64.72 60.75 51.57 20.04
gte-Qwen2-1.5B-instruct 73.90 72.86 72.03 70.77 67.43 65.55 63.47 64.09 60.13 53.86 20.04
e5-mistral-7b-instruct 69.31 67.01 67.64 64.09 57.83 55.53 56.16 51.98 50.10 48.64 20.67
GritLM-7B 79.75 82.25 76.62 78.08 77.66 75.16 75.16 77.87 78.91 78.29 1.46
LLM2Vec 84.34 81.21 79.12 83.09 80.58 79.75 78.71 79.33 80.79 78.50 5.85
Reranker
bge-reranker-v2-m3 81.21 91.02 86.64 86.43 79.33 75.37 75.78 66.39 72.65 62.00 19.21
bge-reranker-v2-gemma 89.35 92.48 90.19 92.90 91.02 85.80 85.39 82.67 80.79 70.35 19.00
followIR 92.28 90.81 89.35 89.56 86.22 84.97 81.42 86.22 76.41 71.82 20.46
Table 16: Win Rate Score as the Number of Conditions in the Query Increases on the Medical Case Dataset (Task 1)
Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 decline
Spare Retriever
BM25 19.72 34.04 30.05 36.85 40.85 41.78 43.66 43.90 43.19 49.77 -30.05
Dense Retriever
jina-embeddings-v3 69.01 63.62 68.31 65.96 62.68 63.85 63.38 59.86 59.86 52.58 16.43
gte-large-en-v1.5 77.70 71.36 70.42 61.03 69.95 64.32 64.32 59.86 57.75 50.23 27.46
NV-Embed-v2 84.27 83.57 77.70 71.83 75.35 70.89 65.73 60.33 58.69 51.17 33.10
bge-en-icl 88.50 78.40 73.47 70.42 68.78 65.73 65.96 61.97 52.35 49.53 38.97
gte-Qwen2-7B-instruct 77.46 72.54 68.31 64.55 60.33 56.57 54.69 52.35 46.24 41.31 36.15
gte-Qwen2-1.5B-instruct 81.46 75.12 72.54 70.19 69.72 66.90 62.21 55.16 56.10 49.30 32.16
e5-mistral-7b-instruct 83.10 71.83 65.02 66.90 61.97 59.86 49.77 51.17 46.71 44.13 38.97
GritLM-7B 85.68 76.53 73.24 67.61 67.14 67.84 64.55 62.21 62.91 64.55 21.13
LLM2Vec 84.98 71.60 67.61 64.08 57.98 58.69 55.16 52.35 46.01 37.79 47.18
Reranker
bge-reranker-v2-m3 88.50 80.99 71.36 66.43 66.43 56.10 54.46 42.25 25.59 13.62 74.88
bge-reranker-v2-gemma 93.43 88.26 82.86 72.77 72.07 69.72 60.56 55.40 43.66 27.23 66.20
followIR 61.74 62.68 54.46 54.46 53.99 48.36 36.85 17.61 9.86 4.93 56.81
Table 17: Win Rate Score as the Number of Conditions in the Query Increases on the Legal Documents Dataset (Task 1)
D.2 Complete Results Of Task 2
Table 18 presents the experimental results of Task 2, where Win Rate reflects the success rate between documents that satisfy different numbers of conditions under a multi-condition query (query10, which contains ten conditions), i.e., $d_{k}$ vs. $d_{k-1}$.
Model $d_{1}$_vs_$d_{0}$ $d_{2}$_vs_$d_{1}$ $d_{3}$_vs_$d_{2}$ $d_{4}$_vs_$d_{3}$ $d_{5}$_vs_$d_{4}$ $d_{6}$_vs_$d_{5}$ $d_{7}$_vs_$d_{6}$ $d_{8}$_vs_$d_{7}$ $d_{9}$_vs_$d_{8}$ $d_{10}$_vs_$d_{9}$ Avg.
Spare Retriever
BM25 13.91 16.50 16.81 18.14 22.10 29.04 37.78 38.87 39.93 40.19 25.90
Dense Retriever
jina-embeddings-v3 73.43 70.52 67.45 66.66 65.32 63.40 63.15 62.82 65.13 60.35 65.82
gte-large-en-v1.5 76.74 73.85 72.70 69.91 70.32 68.05 67.39 64.09 65.14 62.58 69.08
NV-Embed-v2 82.57 76.39 74.45 72.10 73.27 69.15 69.48 66.74 68.75 71.57 72.45
bge-en-icl 79.40 70.58 69.36 68.13 64.80 63.12 63.01 61.72 61.31 63.69 66.51
gte-Qwen2-7B-instruct 79.84 74.02 70.57 69.97 65.44 60.54 61.35 59.42 59.55 60.40 66.11
gte-Qwen2-1.5B-instruct 74.30 71.80 72.28 68.49 69.32 65.69 67.08 64.97 63.46 65.09 68.25
e5-mistral-7b-instruct 75.11 67.88 62.73 58.61 56.87 54.52 55.26 54.03 56.68 61.94 60.36
GritLM-7B 79.59 77.73 73.40 74.71 75.56 72.15 73.52 71.87 72.01 75.21 74.58
LLM2Vec 83.50 74.25 73.43 72.24 70.36 67.21 66.99 67.07 66.49 67.48 70.90
Reranker
bge-reranker-v2-m3 76.08 68.06 63.83 62.06 60.65 58.35 54.79 50.60 48.57 44.96 58.80
bge-reranker-v2-gemma 87.98 82.08 78.07 77.10 76.21 72.13 68.84 65.63 62.32 56.13 72.65
followIR 61.99 59.82 60.87 60.76 59.91 56.41 51.63 47.93 44.71 43.52 54.76
Table 18: Average Win Rate Comparison Between Documents in Task 2
D.3 Complete Results Of Document Length
Table 19 and Table 20 present the effect of document length on retrieval performance, with documents padded to 512 and 1024 words, respectively. We use repeated filler text, such as “The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again.”, following the setting in Wang et al. (2023). The filler text is inserted between the original document sentences until the total text length reaches 512 or 1024 words.
Model $d_{1}$_vs_$d_{0}$ $d_{2}$_vs_$d_{1}$ $d_{3}$_vs_$d_{2}$ $d_{4}$_vs_$d_{3}$ $d_{5}$_vs_$d_{4}$ $d_{6}$_vs_$d_{5}$ $d_{7}$_vs_$d_{6}$ $d_{8}$_vs_$d_{7}$ $d_{9}$_vs_$d_{8}$ $d^{+}$_vs_$d_{9}$ Avg.
Spare Retriever
BM25 11.91 14.34 14.75 14.45 15.99 24.75 36.58 37.68 38.02 39.34 24.78
Dense Retriever
jina-embeddings-v3 64.8 60.11 58.61 58.31 57.83 56.36 56.42 57.20 58.56 58.59 58.68
gte-large-en-v1.5 66.63 61.15 57.66 59.44 56.26 55.10 53.64 53.14 51.34 54.51 56.89
NV-Embed-v2 68.83 62.87 60.97 62.16 61.78 61.80 62.04 61.10 62.75 64.78 62.91
bge-en-icl 70.55 62.18 59.35 60.33 59.70 60.19 59.28 58.95 60.16 60.23 61.09
gte-Qwen2-7B-instruct 68.63 64.87 61.91 60.28 61.98 58.70 59.28 59.06 60.28 59.69 61.47
gte-Qwen2-1.5B-instruct 69.51 66.38 63.47 61.89 60.13 58.76 58.57 57.36 58.62 62.11 61.68
e5-mistral-7b-instruct 69.98 63.50 60.30 59.3 56.77 54.52 53.05 53.14 51.47 53.99 57.60
GritLM-7B 73.46 70.37 69.19 70.36 68.36 67.05 68.34 61.55 58.38 59.57 66.66
LLM2Vec 70.57 68.37 67.10 67.10 66.35 58.73 36.04 36.81 35.37 37.03 54.35
Reranker
bge-reranker-v2-m3 73.65 66.55 63.28 61.65 54.60 38.54 28.42 27.92 28.14 28.02 47.08
bge-reranker-v2-gemma 82.39 77.63 71.80 70.00 65.10 56.91 41.75 32.01 28.84 29.96 55.64
followIR 50.48 51.78 49.64 46.27 37.60 25.22 24.41 24.51 25.67 24.71 36.03
Table 19: Effect of document length on retrieval performance (padded to 512 words).
Model $d_{1}$_vs_$d_{0}$ $d_{2}$_vs_$d_{1}$ $d_{3}$_vs_$d_{2}$ $d_{4}$_vs_$d_{3}$ $d_{5}$_vs_$d_{4}$ $d_{6}$_vs_$d_{5}$ $d_{7}$_vs_$d_{6}$ $d_{8}$_vs_$d_{7}$ $d_{9}$_vs_$d_{8}$ $d^{+}$_vs_$d_{9}$ Avg.
Spare Retriever
BM25 12.25 14.32 14.86 14.81 15.97 24.48 36.00 37.53 38.85 39.53 24.86
Dense Retriever
jina-embeddings-v3 64.56 59.74 58.13 57.20 55.47 55.90 55.33 53.70 54.75 54.81 56.96
gte-large-en-v1.5 68.62 61.61 58.47 54.77 54.97 54.52 54.44 49.88 39.09 38.34 53.47
NV-Embed-v2 59.23 61.26 62.58 62.81 64.57 63.95 62.98 63.49 65.65 67.55 63.41
bge-en-icl 66.08 61.93 60.83 59.63 59.34 61.04 61.08 51.67 36.00 35.95 55.36
gte-Qwen2-7B-instruct 66.06 63.23 63.36 61.35 59.63 58.56 56.62 57.97 36.61 35.96 55.94
gte-Qwen2-1.5B-instruct 68.46 63.66 62.02 62.21 59.47 60.78 59.38 58.71 35.01 34.91 56.46
e5-mistral-7b-instruct 66.97 61.11 59.05 54.53 54.40 54.47 53.02 54.55 53.88 53.57 56.56
GritLM-7B 71.47 67.97 69.41 56.02 54.49 52.58 54.21 56.35 54.87 57.26 59.46
LLM2Vec 74.81 72.05 71.55 26.85 26.68 26.35 25.24 26.16 26.04 27.27 40.30
Reranker
bge-reranker-v2-m3 76.56 70.75 56.83 18.33 19.86 19.09 18.55 20.75 19.88 21.78 34.24
bge-reranker-v2-gemma 83.48 77.02 66.34 19.76 20.54 19.81 20.64 19.34 20.65 21.46 36.90
followIR 52.36 51.43 19.70 18.35 17.15 17.11 17.78 16.94 17.77 18.75 24.73
Table 20: Effect of document length on retrieval performance (padded to 1024 words).
Appendix E Findings In Constructing MultiCon Dataset
E.1 The Use Of LLM-generated Data in Retrieval
In recent years, artificial datasets generated by LLMs have become a common practice for training and evaluating retrieval models (Su et al., 2024; Lee et al., 2024; Weller et al., 2024a). For instance, E5-Mistral (Wang et al., 2024a) rely entirely on LLM-generated datasets for fine-tuning. While this approach can significantly expand training corpora, prior studies have highlighted its potential drawbacks, including introducing inherit linguistic biases of the underlying LLMs (Shumailov et al., 2024), potentially constraining the retrieval model’s performance and generalizability. Furthermore, purely artificial data often lacks the contextual richness and complexity found in real-world retrieval scenarios (Li et al., 2023c; Wang et al., 2024b), making it difficult to capture the actual needs of users’ queries accurately.
During our dataset construction, we observed similar issues. When using LLM-generated transformations to modify positive documents into hard negatives, the model often restructured expressions to fit its learned patterns, even when explicitly instructed to modify only a few condition-related words while keeping the rest unchanged. For example, in the legal documents dataset, a positive sentence like: “The defendant was convicted of fraud under Section 420 of the Penal Code and sentenced to five years in prison.” was frequently modified by the LLM into a generic pattern, such as: “The defendant was found guilty of fraud and received a prison sentence.”
Similarly, in medical case documents, a sentence like: “The patient reported experiencing persistent chest pain and shortness of breath, leading to a diagnosis of angina.” was often transformed into a standardized version: “The patient was diagnosed with a heart condition after reporting chest pain.”
These modifications eroded the diversity and long-tail characteristics of real-world data, reducing the fine-grained variability necessary for retrieval tasks. Instead of preserving rich domain-specific details, LLM-generated transformations tended to normalize distinct cases into overly generic patterns, which could misrepresent real-world retrieval challenges.
Empirical results further confirm the limitations of fully LLM-generated training data. The E5-Mistral model, which relies entirely on synthetic data, performs the worst on MultiConIR. In Task 1, as shown in Table 1, it exhibits the highest performance decline (16.93%) among retrieval models, and in Task 2, as shown in Table 18, its average win rate (60.36%) is the lowest among retrieval models, trailing the second-worst model (Jina-Embeddings-V2) by 5%. These results reinforce the generalization challenges posed by fully synthetic datasets in retrieval tasks, highlighting the importance of incorporating real-world document structures and constraints in training data.
To mitigate this, our pipeline minimizes document-wide modifications, instead restricting LLM interventions to condition sentences only. This targeted approach preserves real-world data authenticity while introducing controlled semantic perturbations, ensuring that retrieval models are trained on meaningful and realistic hard negatives rather than fully synthetic documents.
E.2 Impact of Different Hard Negative Construction Strategies
To systematically examine the impact of hard negative sentence (HNS) construction on retrieval models, we experimented with two distinct approaches: (1) Key Information Modification – altering critical details while maintaining overall sentence structure (applied to books, movies, medical cases, and legal documents). (2) Keyword Retention with Dummy Information – keeping all original keywords intact while injecting irrelevant dummy information (used for the people dataset).
A key objective of this study was to investigate how different HNS construction strategies affect retrieval difficulty. Our initial hypothesis was that the second approach (retaining keywords but adding dummy information) would pose a greater challenge for retrieval models, particularly Dense Retrievers, since hard negatives in this setting contain all the key terms present in positive documents.
However, our experimental results contradicted this expectation. As shown in Fig.8 on the people dataset, Dense Retrieval models remained highly stable, demonstrating a strong ability to differentiate semantic nuances even when all keywords were retained. This suggests that Dense Retrieval primarily relies on contextual embeddings rather than simple keyword matching, allowing it to distinguish between truly relevant documents and distractors with superficial lexical overlap.
In contrast, Reranker models exhibited a significant performance drop when dealing with dummy-information-based HNS. This suggests that Rerankers are more sensitive to this type of negative construction, likely due to their cross-encoder or generative architectures, which process both the query and document jointly. Since Rerankers typically assign scores based on fine-grained textual relevance, the presence of keyword overlap without genuine semantic alignment may mislead them more than Dense Retrieval models.
These findings highlight important considerations for hard negative sampling in multi-condition retrieval. While Dense Retrievers appear robust to surface-level keyword retention, Rerankers are more vulnerable to semantically misleading negatives, suggesting that future retrieval pipelines should adapt negative sampling strategies based on the target retrieval model architecture.
(a) Task1 on Peole dataset
(b) Task1 on Legal dataset
Figure 8: Impact of different HNS construction strategies.

### v3

@@
 Abstract
-In this paper, we introduce MultiConIR, the first benchmark designed to evaluate retrieval models in multi-condition scenarios.
-Unlike existing datasets that primarily focus on single-condition queries from search engines, MultiConIR captures real-world complexity by incorporating five diverse domains: books, movies, people, medical cases, and legal documents.
-We propose three tasks to systematically assess retrieval and reranking models on multi-condition robustness, monotonic relevance ranking, and query format sensitivity.
-Our findings reveal that existing retrieval and reranking models struggle with multi-condition retrieval, with rerankers suffering severe performance degradation as query complexity increases.
-We further investigate the performance gap between retrieval and reranking models, exploring potential reasons for these discrepancies, and analysis the impact of different pooling strategies on condition placement sensitivity.
-Finally, we highlight the strengths of GritLM and Nv-Embed, which demonstrate enhanced adaptability to multi-condition queries, offering insights for future retrieval models.
+Multi-condition information retrieval (IR) presents a significant, yet underexplored challenge for existing systems.
+This paper introduces MultiConIR, a benchmark specifically designed to evaluate retrieval and reranking models under nuanced multi-condition query scenarios across five diverse domains.
+We systematically assess model capabilities through three critical tasks: complexity robustness, relevance monotonicity, and query format sensitivity.
+Our extensive experiments on 15 models reveal a critical vulnerability: most retrievers and rerankers exhibit severe performance degradation as query complexity increases.
+Key deficiencies include widespread failure to maintain relevance monotonicity, and high sensitivity to query style and condition placement.
+The superior performance of GPT-4o reveals the performance gap between IR systems and advanced LLM for handling sophisticated natural language queries.
+Furthermore, this work delves into the factors contributing to reranker performance deterioration and examines how condition positioning within queries affects similarity assessment, providing crucial insights for advancing IR systems towards complex search scenarios.
 The code and datasets are available at https://github.com/EIT-NLP/MultiConIR
 MultiConIR: Towards Multi-Condition Information Retrieval
-Xuan Lu1,2, Sifan Liu2, Bochao Yin2, Yongqi Li3, Xinghao Chen2,3, Hui Su4, Yaohui Jin1, Wenjun Zeng2, Xiaoyu Shen2thanks: Corresponding author, 1Shanghai Jiao Tong University 2Ningbo Key Laboratory of Spatial Intelligence and Digital Derivative, Institute of Digital Twin, EIT 3Department of Computing, The Hong Kong Polytechnic University 4Meituan Inc.
-xyshen@eitech.edu.cn
+Xuan Lu1,2,  Sifan Liu2thanks: Equal contribution,  Bochao Yin2footnotemark: ,  Yongqi Li3,  Xinghao Chen2,3, Hui Su4,  Yaohui Jin1,  Wenjun Zeng2,  Xiaoyu Shen2thanks: Corresponding author 1Shanghai Jiao Tong University 2Ningbo Key Laboratory of Spatial Intelligence and Digital Derivative, Institute of Digital Twin, EIT 3Department of Computing, The Hong Kong Polytechnic University 4Meituan Inc.
+lux1997@sjtu.edu.cn xyshen@eitech.edu.cn
 1 Introduction
@@
 Traditionally, IR systems retrieve documents by matching queries based on lexical similarity, such as BM25 (Carpineto and Romano, 2012; Ponte and Croft, 2017), or semantic similarity using dense vector representations (Karpukhin et al., 2020; Zhan et al., 2021).
-Though highly effective for queries with straightforward query-document relationships (Su et al., 2024), they often fail to fully capture nuanced intent as user needs become more complex (Zhu et al., 2023; Su et al., 2024).
+Though highly effective for queries with straightforward query-document relationships (Su et al., 2024), they often fail to fully capture nuanced intent as user needs become more complex (Zhu et al., 2023; Su et al., 2024; Zhu et al., 2025).
 Figure 1: From single-condition to multi-condition retrieval.
-Standard retrieval relies on single-condition queries, while instruction-following retrieval enhances queries with explicit instructions.
-Traditional SQL-based filtering handles predefined attributes, whereas real-world multi-condition retrieval involves queries specifying multiple conditions, which can be attribute-level or semantic-level.
-A significant challenge arises when users specify multiple requirements simultaneously, as illustrated in Fig 1.
+Standard and instruction-aware retrieval address single-condition queries.
+SQL-based filtering is restricted to predefined attributes within structured data.
+Real-world multi-condition retrieval enables the formulation of multiple, often semantic, conditions using natural language query.
+A significant challenge arises when users specify multiple requirements simultaneously, as illustrated in Fig.
+1.
 Whether searching for a movie with specific attributes or selecting a product that meets various criteria, multi-condition search has become an integral part of modern information-seeking behavior.
@@
 As a result, it struggles to support nuanced or semantic-level queries that go beyond structured data filtering.
-The advent of Large Language Models (LLMs) has enhanced IR by introducing instruction-following capabilities (Asai et al., 2023; Weller et al., 2024a; Oh et al., 2024).
+The advent of Large Language Models (LLMs) has enhanced IR by introducing instruction-following capabilities (Asai et al., 2023; Weller et al., 2024a; Oh et al., 2024; Liu et al., 2024).
 This approach augments standard queries with explicit instructions, which serve as additional constraints to refine search results, as shown in Fig.1.
 Despite these advancements, existing evaluation benchmarks remain predominantly focused on single-condition queries and binary relevance assessments—classifying documents as either relevant or irrelevant (Nguyen et al., 2016; Kwiatkowski et al., 2019; Muennighoff et al., 2022)—thus overlooking the nuanced challenges of multi-condition queries, where relevance depends on the degree to which multiple conditions are satisfied.
-An ideal multi-condition retrieval system should exhibit the following properties: (1) Complexity Robustness: The system should maintain high performance regardless of query complexity (i.e., the number of conditions specified); (2) Relevance Monotonicity: The relevance scores should scale monotonically with the number of matched conditions; for example, a document matching all $n$ conditions should be ranked higher than one matching $n-1$; (3) Format Invariance: The system should yield consistent results regardless of the query format, whether presented as a structured list or as free-form natural language.
+An ideal multi-condition retrieval system should exhibit the following properties: (1) Complexity Robustness: The system should maintain high performance regardless of query complexity (i.e., the number of conditions specified); (2) Relevance Monotonicity: The relevance scores should scale monotonically with the number of matched conditions; for example, a document matching all $n$ conditions should be ranked higher than one matching n1n-1; (3) Format Invariance: The system should yield consistent results regardless of the query format, whether presented as a structured list or as free-form natural language.
 Existing benchmarks do not offer a structured framework for evaluating multi-condition retrieval along these dimensions.
-To address this gap, we introduce MultiCondIR—the first benchmark designed to comprehensively evaluate multi-condition retrieval systems.
-Through systematic experiments on 12 state-of-the-art models (including dense retrievers, cross-encoders, and LLM-based agents), we uncover several critical insights:
-  • Multi-Condition Degradation: Dense retrieval and reranking models deteriorate as query conditions increase, with rerankers suffering more severely.
-Rerankers’ token-level interactions benefit single-condition tasks but hinder multi-condition retrieval.
-  • Relevance Monotonicity Failure: Models fail to maintain a consistent ranking hierarchy.
-Dense retrievers tend to focus more on “exact match” and “complete mismatch”, likely due to their training paradigm.
-  • Query Format Sensitivity: Models are susceptible to query phrasing, with rerankers exhibiting a flip rate exceeding 20%.
-  • Pooling Impact on Focus: Different embedding pooling strategies influence the model’s focus on query condition positions.
-Mean pooling focuses on early query conditions, <EOS> pooling biases toward later conditions, while latent attention layer pooling distributes focus more evenly.
-  • GritLM’s Advantage: GritLM outperforms other models, potentially due to its hybrid attention mechanism, which integrates both bidirectional and causal attention to improve the understanding of complex queries.
-By quantifying these gaps, our work provides actionable guidelines for developing IR systems that truly understand multi-condition intent, laying the groundwork for advancing IR toward human-like reasoning in complex search scenarios.
+To address this gap, we introduce MultiConIR—a benchmark designed to comprehensively evaluate multi-condition retrieval systems.
+Through systematic experiments on 15 state-of-the-art models (including dense retrievers, cross-encoders, and LLM-based agents), we uncover several critical insights:
+  • Multi-Condition Struggle: Retrievers and Rerankers struggle with multi-condition retrieval, showing performance decline as query conditions increase, difficulty with relevance monotonicity, and sensitivity to query style variations.
+  • Retrievers and Rerankers Differ: Rerankers excel with single-condition queries but fail under multiple conditions.
+Retrievers demonstrate greater robustness.
+GritLM demonstrates the best robustness among retrievers.
+  • Position Impacts Model Focus: Dense retriever pooling strategies emphasize different condition positions, mean pooling focuses on initial positions, while <EOS> pooling emphasizes final positions.
+Rerankers exhibit non-uniform attention across positions and greater sensitivity to document length variations.
+By quantifying these gaps, our work reveals key deficiencies in the ability of current IR systems to understand multi-condition intent, laying the groundwork for advancing IR toward human-like reasoning in complex search scenarios.
 2 Related Works
@@
 Recent advancements have shown that LLMs offer significant potential as backbone encoders for dense retrieval (Wang et al., 2024a; Weller et al., 2024c; BehnamGhader et al., 2024).
-For instance, Repllama (Ma et al., 2023) enhanced retrieval performance by fine-tuning Llama-2 to serve as dense retrievers.
+For instance, Repllama (Ma et al., 2023) fine-tuned Llama-2 to serve as dense retrievers.
 GritLM (Muennighoff et al., 2024) unified text embedding and generation within a single LLM.
@@
 Recent studies have expanded retrieval benchmarks to address more complex scenarios.
+BIRCO (Wang et al., 2024b) introduces a benchmark designed to evaluate information retrieval tasks with complex objectives.
 Instruction-based datasets (Weller et al., 2024a; Qin et al., 2024; Oh et al., 2024), for instance, evaluate the instruction-following capabilities of retrieval models by embedding explicit instructions within queries to better represent users’ retrieval intents.
@@
 (2024b), and multi-hop reasoning Su et al.
+(2024); Liu et al.
 (2025).
@@
 (2) A positive document $d^{+}$ that satisfies all $k$ conditions, i.e., $d^{+}\models C$,
-(3) A sequence of hard negative documents $\{d_{0},d_{1},\dots,d_{k-1}\}$, where each $d_{j}$ satisfies exactly $j$ out of $k$ conditions, formally expressed as $d_{j}\models\{c_{1},\dots,c_{j}\}$ and $d_{j}\not\models\{c_{j+1},\dots,c_{k}\}$.
+(3) A sequence of hard negative (HN) documents $\{d_{0},d_{1},\dots,d_{k-1}\}$, where each $d_{j}$ satisfies exactly $j$ out of $k$ conditions, formally expressed as $d_{j}\models\{c_{1},\dots,c_{j}\}$ and $d_{j}\not\models\{c_{j+1},\dots,c_{k}\}$.
 This controlled design enables a principled evaluation of multi-condition retrieval along three fundamental axes: (1) Complexity Robustness: The model’s retrieval effectiveness as $k$ increases, measured by its ability to distinguish $d^{+}$ from $d_{k-1}$; (2) Relevance Monotonicity: The extent to which the retrieval model enforces a strict ordering such that $S(q_{k},d_{j})>S(q_{k},d_{j+1})$ for all $j$, ensuring that documents satisfying more conditions are ranked higher; and (3) Format Invariance: The stability of retrieval performance under transformations of query representation, quantified by discrepancies in ranking outcomes across query formats.
 3.1 Domain Selection
-MultiConIR was constructed from real-world datasets from five diverse domains: books, movies, people, medical cases, and legal documents.
-These domains were carefully chosen for their practical significance and their intrinsic requirement for multi-condition retrieval.
-Books & Movies: Queries in these domains often blend attribute-based (e.g., a sci-fi movie) with narrative-based (e.g., a story about time travel), requiring semantic understanding beyond keyword matching.
-People: Queries are often based on vague recollections or specific but incomplete information about a person (e.g., a Nobel laureate in Physics who studied black holes) Medical Case & Legal Document: These domains require fine-grained condition matching, where case-based reasoning is essential (e.g., a patient with breathing issues and antibiotic allergies), demanding models understand condition dependencies, not just match isolated terms.
+To construct the MultiConIR dataset, we meticulously selected five domains—Books, Movies, People, Medical Cases, and Legal Documents—each chosen for its practical significance and inherent suitability for evaluating multi-condition retrieval capabilities.
+Books & Movies: These domains represent common consumer searches where nuanced preferences are expressed by combining structured attributes (e.g., genre, creator, year, cast) with narrative elements (e.g., plot details, thematic content like “a story about time travel”).
+Effective retrieval demands semantic understanding beyond simple keyword matching to process multifaceted queries, such as "an action film directed by Christopher Nolan, starring Leonardo DiCaprio, released after 2010, with an intense chase scene."
+People: Queries about individuals frequently rely on partial or vague information, such as notable achievements or specific traits.
+An example query could be “a Nobel laureate in Physics who studied black holes.” These searches demand that IR systems effectively handle incomplete data and infer connections between various attributes to identify the correct individual.
+Medical Case & Legal Document: The medical case and legal document domains offer more practical and application-driven use cases.
+In the medical domain, doctors often rely on retrieval systems to reference historical cases to support diagnostic decisions.
+A typical query might include multiple conditions, such as “Find a case that meets the following conditions: 1) middle-aged female patient; 2) hospitalized for breathing difficulties; 3) has a history of antibiotic allergies; 4) has a family history of peanut allergies.” Similarly, in the legal domain, retrieval users often seek case law with high similarity to ongoing cases, which requires matching various legal and factual attributes in historical court decisions.
+These complex queries require IR systems to perform fine-grained condition matching and understand the interdependencies between various factors.
 3.2 Dataset Construction Pipeline
-Figure 2: MultiCon Dataset Construction Pipeline.
-To construct MultiConIR, we design a multi-step data generation framework as shown in Fig.
+Figure 2: MultiConIR Dataset Construction Pipeline: (1) relevant condition sentences are extracted from documents; (2) these conditions are then used to generate multi-condition queries; (3) hard negative (HN) versions of the condition sentences are created; and (4) positive documents and progressively challenging HN documents are assembled from these elements.
+To construct MultiConIR, we design a multi-step data generation framework.
+As shown in Fig.
 2, this pipeline is highly adaptable across multiple domains, enabling the generation of queries and hard negative (HN) documents that progressively satisfy 1 to 10 conditions.
-To preserve dataset integrity and mitigate the generalization issues associated with fully synthetic datasets (Li et al., 2023c; Wang et al., 2024b), we employ LLM-based generation (GPT-4o) exclusively for modifying sentences within hard negatives, rather than altering entire documents.
-[^1] Discussion about the challenges associated with fully LLM-generated datasets can be found in Appendix E.1.
+To preserve dataset integrity and mitigate the generalization issues associated with fully synthetic datasets (Li et al., 2023c; Wang et al., 2024c), we employ LLM-based generation (GPT-4o) exclusively for modifying sentences within hard negatives, rather than altering entire documents.
+[^1] Fully LLM-generated datasets may introducing inherent linguistic biases of the underlying LLMs, and lacks the contextual richness and complexity in real-world retrieval (Shumailov et al., 2024).
+To mitigate these issues, we restricted LLM interventions to modifying only condition sentences rather than entire document.
+We further discuss this problem in Appendix G.1.
 The data creation process consists of the following steps, with detailed prompt templates provided in Appendix B:
-Step 1: Condition Sentence Extraction.
-To capture the fine-grained constraints for multi-condition retrieval, we prompt GPT-4o to extract ten key sentences from each document, where each sentence represents a distinct and semantically complete condition.
-To ensure robustness, extracted sentences must be contextually relevant, non-redundant, and representative of the document’s key information.
+Step 1: Condition Sentences Extraction.
+For each source document $d$, we issue a structured prompt to GPT-4o that (1) reads the entire text, (2) identifies ten non-overlapping conditions that describe key features expressed in $d$, and (3) return ten original sentences, where every sentence is semantically complete, and expresses a distinct condition.
+Documents yielding fewer than ten qualified sentences are discarded.
+The final ten sentences constitute the condition set $C(d)$, which we later recombine into queries of varying complexity and use as the ground-truth positive for multi-condition retrieval experiments.
 Step 2: Query Generation.
-Given the extracted condition sentences, we prompt GPT-4o to generate ten hierarchical queries (Query1 to Query10), each incorporating an increasing number of conditions (from 1 to 10).
-To enhance linguistic diversity, queries are formulated in two distinct styles: (1) Instruction-style (explicitly listing conditions in a structured format).
-(2) Descriptive-style (embedding conditions naturally within a coherent sentence).
+Given the ten-sentence condition set $C(d)$, we prompt GPT-4o to synthesise a hierarchy of ten queries $\{q_{1},q_{2},...,q_{10}\}$.
+Each $q_{k}$ incorporating the first $k$ conditions.
+To enhance linguistic diversity, For every $q_{k}$ we request two forms: (1) Instruction-style $q_{k}^{\text{inst}}$: a bullet-like template (“Find a document that satisfies: (1)… (2)…”) for explicit parsing.
+(2) Descriptive-style $q_{k}^{\text{des}}$: embedding conditions naturally within a coherent sentence.
 Step 3: Hard Negative Sentence Construction.
-To create hard negatives, we prompt GPT-4o to modify extracted condition sentences into Hard Negative Sentences (HNS) that subtly deviate from the original meaning while preserving lexical plausibility.
-Unlike naive negation, these modifications introduce semantic shifts that retain fluency but fail to satisfy the original retrieval conditions, thereby making retrieval harder.
-We experimented with two approaches for HNS construction: one modifies key information (applied to books, movies, medical cases, and legal documents), while the other retains keywords but adds dummy information (used for the people dataset).
-[^2] The differences between these two approaches are discussed in Appendix E.2.
-Step 4: HN Document Generation.
-To construct multi-condition retrieval datasets, we systematically assemble documents: (1) Positive documents: Composed of all ten original condition sentences, ensuring full query relevance.
-(2) Hard Negative documents $d_{1}$ to $d_{10}$: Generated by progressively replacing 1 to 10 original condition sentences with their corresponding hard negative versions, creating a continuum of relevance degradation.
-This structured data generation ensures that each query condition incrementally refines document ranking, allowing for fine-grained retrieval performance assessment.
+For each condition sentence $c_{i}\in C(d)$ we instruct GPT-4o to produce a semantically divergent yet fluently written variant $h_{i}$ that no longer satisfies the original constraint.
+The rewrite must preserve overall length and style while introducing either (i) a subtle alteration of a critical fact (applied to books, movies, medical cases, and legal documents) or (ii) an innocuous clause that injects misleading information without changing the existing keywords (used for the people corpus).
+[^2] Retrievers are more robust when adding misleading information; Modifying critical facts is challenging for both retrievers and rerankers.
+A detailed comparison of these two strategies is given in Appendix G.2.
+The ten variants form the hard-negative set $\mathrm{HNS}(d)=\{h_{1},\dots,h_{10}\}$.
+Step 4: Hard Negative Document Generation.
+Starting from the ten-sentence condition set $C(d)=\{c_{1},\dots,c_{10}\}$ and its hard-negative counterparts $\mathrm{HNS}(d)=\{h_{1},\dots,h_{10}\}$, we build an ordinal ladder of document variants.
+The positive document $d^{+}$ contains the full sequence $c_{1},\dots,c_{10}]$.
+For each $k\in\{1,\dots,10\}$ we generate a HN document $d_{k}^{-}=\bigl{[}h_{1},\dots,h_{k},\;c_{k+1},\dots,c_{10}\bigr{]},$ i.e., the first $k$ conditions are replaced by their semantically perturbed versions while the remaining 10k10-k conditions stay intact.
+This yields a controlled degradation chain $d^{+}\!\rightarrow d_{1}^{-}\!\rightarrow\cdots\!\rightarrow d_{10}^{-}$, ranging from a single violated constraint to a completely adversarial variant.
+Coupled with the hierarchical queries $\{Q_{k}\}_{k=1}^{10}$, the corpus enables fine-grained evaluation of retrieval performance under progressively stricter condition sets.
+Figure 3: Benchmark quality evaluation framework of MultiConIR.
+Query realism was assessed by human annotators.
+Label accuracy involved initial GPT-4o filtering, followed by a final human double-check.
+Benchmark Quality Assurance.
+The reliability of MultiConIR was audited on two complementary fronts (Fig.
+3).
+(1) Query validity.
+From each of the five domains we randomly sampled 100 multi-condition queries (500 in total) and asked ten trained annotators to judge whether each query was natural, precise, and contextual plausibility (valid vs.
+unvalid).
+The resulting inter-annotator agreement reached $93.7\%$, with Fleiss’ $\kappa=0.84$.
+(2) Document–label validity.
+To detect false positives/negatives, we first applied GPT-4o to the entire corpus: the model verified for every document variant $d_{k}^{-}$ whether exactly $k$ of the ten conditions were satisfied, and we discarded mismatched instances.
+We then drew another 100 document–query pairs per domain (500 total) for manual spot-checking; two independent annotators reviewed each pair and a third adjudicated disagreements, yielding a residual error rate of $2.4\%$.
+After these filtering steps the final benchmark sizes for each domain are summarised in Table 1, and the full evaluation protocol is detailed in Appendix C.
+Domain Number Avg. Length Source License Example
+$D^{+}$ $Q$ $D^{-}$ $Q$ $D$
+People 420 4200 4200 31.2 225.2 People Wikipedia Dat(Mahajan, 2017) CC0 1.0 Table 8
+Books 482 4820 4820 25.6 235.3 Books Dataset (Rustamov, 2021) CC0 1.0 Table 9
+Movies 500 5000 5000 24.8 184.6 Wikipedia Movie Plots (Robischon, 2018) CC BY-SA 4.0 Table 10
+Medical Case 479 4790 4790 28.4 212.1 Medical Transcription (MTSamples, 2025) CC0 1.0 Table 11
+Legal Document 426 4260 4260 34.1 302.2 LexGLUE (Chalkidis et al., 2022) CC BY 4.0 Table 12
+Table 1: Data statistics of MutiConIR.
+For each dataset, we show the number of positive documents ($D^{+}$ ), queries ($Q$) and hard negative documents($D^{-}$), and the average length (in words) of queries and documents, and the source dataset of each domain.
 3.3 Evaluation Metrics
+Conventional IR metrics—e.g., Precision@1, NDCG@kk—only confirm whether a highly relevant document appears early, but they cannot distinguish how well a model orders candidates that satisfy different numbers of query conditions.
+We therefore propose Win Rate: the proportion of pairwise comparisons in which a candidate that fulfils more conditions ranks above one that fulfils fewer.
+We further discuss the difference of Win Rate and traditional IR metrics in the Appendix D
 Complexity Robustness
@@
 The candidate set comprises a Positive document that fully satisfies all conditions and a HN1 document, which is derived from the positive by modifying a single condition.
-Complexity robustness is measured using Win Rate (WR) under various $k$, defined as:
-$\text{WR}_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(S(q_{k},d^{+})>S(q_{k},d_{k-1})),$ (1)
+Complexity robustness is measured using Win Rate (WR)  [^3] In Complexity Robustness evaluation, $\text{WR}_{k}$ and Precision@1 are numerically equivalent.
+under various $k$, defined as:
+$\text{WR}_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(S(q_{k},d^{+})>S(q_{k},d_{k-1})),$
 where $S(q_{i},d^{+})$ and $S(q_{i},d^{-})$ denote similarity scores for the positive document and hard negative.
@@
 The query is fixed as query10 (containing all 10 conditions), while the candidate set includes one positive and ten hard negatives ($d_{0}-d_{9}$), each containing 0–9 conditions.
-We evaluate performance using Win Rate between adjacent hard negatives:
-$\begin{split}\text{WR}_{k,k-1}&=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big{(}S(q_{10},d_{k})\\ &\quad>S(q_{10},d_{k-1})\big{)},\end{split}$ (2)
+We evaluate performance using $\text{WR}_{k,k-1}$ between adjacent hard negatives:
+$\begin{split}\text{WR}_{k,k-1}&=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big{(}S(q_{10},d_{k})\\ &\quad>S(q_{10},d_{k-1})\big{)},\end{split}$
 Format invariance.
@@
 To quantify ranking variability between query styles, we define the Flip Rate (FR):
-$\begin{split}\text{FR}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big{(}\mathbf{1}(S_{\text{inst}}(q_{10},d_{k})>S_{\text{inst}}(q_{10},d_{k-1}))\\ \neq\mathbf{1}(S_{\text{desc}}(q_{10},d_{k})>S_{\text{desc}}(q_{10},d_{k-1}))\big{)},\end{split}$ (3)
+$\begin{split}\text{FR}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big{(}\mathbf{1}(S_{\text{inst}}(q_{10},d_{k})>S_{\text{inst}}(q_{10},d_{k-1}))\\ \neq\mathbf{1}(S_{\text{desc}}(q_{10},d_{k})>S_{\text{desc}}(q_{10},d_{k-1}))\big{)},\end{split}$
 where $S_{\text{inst}}$ and $S_{\text{desc}}$ denote similarity scores under instruction-style and descriptive-style queries.
@@
 A higher FR indicates greater sensitivity to query formulation.
-Figure 3: Effect of Condition Quantity on Average Win Rate (reflecting complexity robustness).
-Win Rate measures Positive vs.
-Hard Negative success as query conditions increase.
-Blue, green and red lines represent Dense Retrievers, Rerankers and Red line respectively.
-Figure 4: Relevance Monotonicity Distinction.
-Win Rate reflects the success rate between documents satisfying different numbers of conditions under a multi-condition query (query10), i.e., $d_{k}$ vs.
-$d_{k-1}$.
 4 Experiments
-We evaluate 13 representative retrieval models from diverse architectures and varying model sizes, including one spare retrieval model: BM25 (Robertson and Zaragoza, 2009); Two BERT-based retrieval models: gte-large-en-v1.5 (Li et al., 2023b) and jina-embeddings-v3 (Sturua et al., 2024); Seven LLM-based retrieval models: NV-Embed-v2 (Lee et al., 2024), bge-en-icl (Li et al., 2024), gte-Qwen2-7B-instruct (Li et al., 2023b), gte-Qwen2-1.5B-instruct (Li et al., 2023b), e5-mistral-7b-instruct (Wang et al., 2024a), GritLM-7B (Muennighoff et al., 2024), LLM2Vec (BehnamGhader et al., 2024); Three reranking models: bge-reranker-v2-m3 (Chen et al., 2024), bge-reranker-v2-gemma (Chen et al., 2024), FollowIR-7B (Weller et al., 2024a).
+We evaluate 15 representative retrieval and reranking models from diverse architectures and varying model sizes, including sparse retrieval model: BM25 (Robertson and Zaragoza, 2009); BERT-based retrieval models: gte-large-en-v1.5 (Li et al., 2023b) and jina-embeddings-v3 (Sturua et al., 2024); LLM-based retrieval models: NV-Embed-v2 (Lee et al., 2024), bge-en-icl (Li et al., 2024), gte-Qwen2-7B-instruct (Li et al., 2023b), gte-Qwen2-1.5B-instruct (Li et al., 2023b), e5-mistral-7b-instruct (Wang et al., 2024a), GritLM-7B (Muennighoff et al., 2024), LLM2Vec (BehnamGhader et al., 2024); Point-wise reranking models: bge-reranker-v2-m3 (Chen et al., 2024), bge-reranker-v2-gemma (Chen et al., 2024), FollowIR-7B (Weller et al., 2024a); Fine-tuned list-wise reranker: RankZephyr (Pradeep et al., 2023); Advanced LLM for zero-shot ranking: GPT-4o (OpenAI, 2024).
 Details of each model are provided in Appendix A.
 Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 Decline
-Spare Retriever
-BM25 28.59 34 33.14 36.53 37.38 37.95 38.08 38.41 38.86 39.87 -11.28
+Sparse Retriever
+BM25 28.59 34 33.14 36.53 37.38 37.95 38.08 38.41 38.86 39.87 $\downarrow$-11.28
 Dense Retriever
-jina-embeddings-v3 76.09 71.26 71.84 71.04 65.59 65.65 64.75 64.24 64.62 60.71 15.38
-gte-large-en-v1.5 75.87 77.26 73.79 70.22 70.71 67.40 67.53 64.97 65.36 61.36 14.51
-NV-Embed-v2 80.53 80.32 78.81 75.70 75.68 72.61 73.28 71.54 70.00 68.02 12.51
-bge-en-icl 83.42 80.65 78.44 76.77 74.54 73.00 74.23 73.25 69.70 68.00 15.42
-gte-Qwen2-7B-instruct 70.75 72.22 69.99 68.51 65.20 63.53 62.22 62.20 59.15 56.17 14.58
-gte-Qwen2-1.5B-instruct 73.64 74.97 72.23 71.37 69.94 67.46 66.92 64.64 63.65 58.68 14.96
-e5-mistral-7b-instruct 75.05 70.85 68.18 67.45 63.70 61.60 59.70 59.07 57.85 58.12 16.93
-GritLM-7B 82.08 80.32 78.38 76.40 76.40 73.50 75.69 74.62 74.53 75.95 6.13
-LLM2Vec 83.13 77.42 75.49 75.48 72.49 72.56 70.56 70.73 68.71 67.00 16.13
-Reranker
-bge-reranker-v2-m3 87.14 85.56 78.62 76.05 74.29 68.41 67.86 59.48 55.59 44.87 42.27
-bge-reranker-v2-gemma 91.07 90.02 86.70 84.99 83.17 79.00 75.89 72.29 67.11 56.09 34.98
-followIR 83.41 79.72 76.25 74.60 70.12 67.94 62.62 55.93 48.59 43.52 39.89
+jina-embeddings-v3 76.09 71.26 71.84 71.04 65.59 65.65 64.75 64.24 64.62 60.71 $\downarrow$15.38
+gte-large-en-v1.5 75.87 77.26 73.79 70.22 70.71 67.40 67.53 64.97 65.36 61.36 $\downarrow$14.51
+NV-Embed-v2 80.53 80.32 78.81 75.70 75.68 72.61 73.28 71.54 70.00 68.02 $\downarrow$12.51
+bge-en-icl 83.42 80.65 78.44 76.77 74.54 73.00 74.23 73.25 69.70 68.00 $\downarrow$15.42
+gte-Qwen2-7B-instruct 70.75 72.22 69.99 68.51 65.20 63.53 62.22 62.20 59.15 56.17 $\downarrow$14.58
+gte-Qwen2-1.5B-instruct 73.64 74.97 72.23 71.37 69.94 67.46 66.92 64.64 63.65 58.68 $\downarrow$14.96
+e5-mistral-7b-instruct 75.05 70.85 68.18 67.45 63.70 61.60 59.70 59.07 57.85 58.12 $\downarrow$16.93
+GritLM-7B 82.08 80.32 78.38 76.40 76.40 73.50 75.69 74.62 74.53 75.95 $\downarrow$6.13
+LLM2Vec 83.13 77.42 75.49 75.48 72.49 72.56 70.56 70.73 68.71 67.00 $\downarrow$16.13
+Fine-tuned Reranker
+bge-reranker-v2-m3 87.14 85.56 78.62 76.05 74.29 68.41 67.86 59.48 55.59 44.87 $\downarrow$42.27
+bge-reranker-v2-gemma 91.07 90.02 86.70 84.99 83.17 79.00 75.89 72.29 67.11 56.09 $\downarrow$34.98
+followIR 83.41 79.72 76.25 74.60 70.12 67.94 62.62 55.93 48.59 43.52 $\downarrow$39.89
+RankZephyr 92.72 90.29 88.38 87.69 84.57 80.88 78.93 75.99 72.39 66.84 $\downarrow$25.88
+Zero-shot LLM for Ranking
+GPT-4o 95.49 94.89 93.71 92.11 90.81 89.08 88.43 88.08 86.82 85.26 $\downarrow$9.23
 Table 2: Impact of increasing condition quantity in queries on average Win Rate (Task 1). The Decline reflects the degree of Win Rate reduction from query1 to query10.
 4.1 Results for Complexity Robustness
-Table 1 presents the average Win Rate scores for evaluating complexity robustness across five datasets [^3] Individual dataset performances are detailed in the Appendix D.1..
-To visualize how model performance evolves as the number of query conditions increases, we plot a line graph with Gaussian smoothing (window size = 1), as shown in Fig.
-4.
-The results reveal several notable trends:
-Performance decline with increasing query conditions
+Table 2 presents the average Win Rate scores for evaluating complexity robustness across five datasets. The results reveal several notable trends:
+Performance decline with more conditions
 As the number of conditions in the query increases, the performance of both retrieval and reranking models declines.
 This suggests that with more conditions, models struggle to accurately distinguish between positives and HNs.
-Among all models, GritLM-7B exhibits the lowest performance degradation, declining by only 6.13% from Query1 to Query10.
-NV-Embed-v2 follows with a relatively low decline of 12.51%.
-Reranking models exhibit steeper performance drop
-Initially, reranking models outperform retrieval models with single-condition queries.
+Among all models, GPT-4o maintained the highest win rate from Query1 to Query10, with its performance declining by 9.23%.
+GritLM-7B exhibits the lowest performance degradation of 6.13%.The remaining models all exceeded 10% decline.
+Rerankers exhibit steeper performance drop
+As shown in Table 2, fine-tuned rerankers outperform retrievers with single-condition queries.
 However, as the number of conditions increases, their performance declines more sharply.
-As shown in Table 1, for Query1, rerankers almost outperformed all retrieval models.
-bge-reranker-v2-gemma achieves a Win Rate of 91.07% at Query1 but drops to 56.09% at Query10, a drastic 34.98% decrease.
-Similarly, bge-reranker-v2-m3 declines from 87.14% to 44.87%, a drop of 42.27%, and followIR shows a drop from 83.41% to 43.52%, with a drop of 39.89%.
-In contrast, retrievers exhibit an average Win Rate decline of 14.06%.
-Sparse retrieval models show performance improvement
-Unlike dense retrievers and rerankers, BM25’s performance improves with query complexity, increasing from 28.59% (Query1) to 39.87% (Query10).
-This is likely due to its lexical matching mechanism, where more query conditions lead to greater lexical overlap with candidate documents, enhancing retrieval effectiveness.
-However, BM25’s overall win rate remains lower than dense retrievers, and positives rank below HNs (win rate < 0.5), suggesting that hard negatives consistently outrank positives.
-This may stem from dataset construction, where queries and HNs were generated or paraphrased by GPT-4o, potentially introducing lexical bias [^4] Appendix E further explores challenges in using generated datasets..
+Eventually, rerankers even fell behind some retrievers.
+The Win Rates for all rerankers declined by over 25%, with an average decline of 35.76%.
+For retriever models, the average decline was 14.06%.
 4.2 Results for Relevance Monotonicity
 Fig.
-4 illustrates the trend of average Win Rate in the multi-condition retrieval setting of Task 2555The complete results are provided in Appendix D.2, which evaluates the model’s ability to distinguish the relevance hierarchy among documents with varying conditions.
+4 illustrates the trend of average $\text{WR}_{k,k-1}$ in the multi-condition retrieval setting of Task 2, which evaluates the model’s ability to distinguish the relevance hierarchy among documents with varying conditions.
+The complete results are provided in Table 13.
 Several key observations can be made:
-Relevance monotonicity failure
-As documents become increasingly hard (i.e., satisfying more conditions in the query), retrieval and reranking models struggle to distinguish between $d_{k}$ and $d_{k-1}$, leading to a decline in Win Rate performance.
+Relevance monotonicity struggle
+As documents become increasingly hard (i.e., satisfying more conditions in the query), it becomes harder for retrieval and reranking models to accurately distinguish $d_{k}$ and $d_{k-1}$, leading to a decline in Win Rate performance.
 This failure emphasizes the challenge of preserving relevance monotonicity in multi-condition retrieval settings and highlights a gap in current model capabilities when handling complex queries.
-Sensitive to “exact match” and “complete mismatches”
+Sensitive to exact match and complete mismatches
 We observe a slight upward trend at the end of Win Rate curves for most dense retrievers, likely due to their contrastive learning-based training.
@@
 4.3 Results for Format Invariance
+Table 3 presents the Flip Rate induced by query format variations.
+GPT-4o showed the lowest flip rate of 6.98%, showcasing the robustness of advanced LLMs against variations in query style.
+Additionally, most models exceed 10%, indicating a substantial impact of query formulation on retrieval performance.
+Table 3 presents the complete results of Format Invariance.
 Model People Books Movies Medical Legal Avg.
-Spare Retriever
+Sparse Retriever
 BM25 14.86 17.88 16.84 19.25 12.14 16.19
@@
 LLM2Vec 12.81 7.93 9.56 8.12 10.49 9.78
-Reranker
+Fine-tuned Reranker
 bge-reranker-v2-m3 42.40 32.22 34.82 28.35 31.24 33.81
@@
 followIR 35.81 31.60 23.70 25.07 28.43 28.92
+RankZephyr 20.21 16.34 14.86 11.53 25.31 17.65
+Zero-shot LLM for Ranking
+GPT-4o 7.21 4.22 6.78 6.32 10.35 6.98
 Table 3: Flip Rate for query format shift (Task 3). The Flip Rate reflects the win rate reversal when switching the query format from instruction-style to descriptive-style.
-Table 2 presents the Flip Rate induced by query format variations.
-Most models exceed 10%, indicating a substantial impact of query formulation on retrieval performance.
-BM25 records a 16.19% Flip Rate, which can serve as a baseline for the lexical differences caused by query format variations.
-Dense retrieval models show relatively lower sensitivity, with Flip Rates between 8% to 16%.
-GritLM-7B (8.21%), NV-Embed-v2 (9.12%), and LLM2Vec (9.78%) exhibit less variation, suggesting higher robustness to query styles.
+Dense retrieval models show relatively lower sensitivity than rerankers, with Flip Rates between 8% to 16%.
+GritLM-7B (8.21%), NV-Embed-v2 (9.12%), and LLM2Vec (9.78%) exhibit less variation.
 In contrast, reranking models show significantly higher sensitivity to query format changes, with Flip Rates exceeding 20%.
-The highest Flip Rate observed is 33.81% for bge-reranker-v2-m3, confirming their susceptibility to query formulation changes.
+The highest Flip Rate observed is 33.81% for bge-reranker-v2-m3.
+Figure 4: Relevance Monotonicity Distinction. Win Rate reflects the success rate between documents satisfying different numbers of conditions under a multi-condition query.
 (a) Retrieval Performance when padding to 512 words
-(b) Retrieval Performance when padding to 512 words
+(b) Retrieval Performance when padding to 1024 words
 Figure 5: Retrieval performance when padding the document set to 512 words and 1024 words.
+Rerankers are highly sensitive to increases in document length, showing rapid performance degradation, whereas retrievers remain comparatively robust.
 5 Analysis
 5.1 Retrievers vs. Rankers
-Our experiments reveal notable differences between retrieval and reranking models across the three tasks.
-Specifically, retrieval models demonstrate greater robustness to query complexity, better preservation of relevance monotonicity, and stronger invariance to query format changes compared to rerankers.
-We analyze these differences through the lens of relevance computation and attention mechanisms.
-Retrieval models: robustness induced by bidirectional attention and dual-encoder.
-Retrieval models typically employ a dual-encoder architecture, where queries and documents are independently encoded before computing their similarity using dot-product or cosine similarity.
-This independent computation ensures that the generation of query and document embeddings remains unaffected by each other.
-At the same time, bidirectional attention enables the model to better capture the overall semantic meaning of the query.
-Another critical factor is the training datasets.
-Many dense retrieval models are trained on a mixture of retrieval-specific and general textual datasets (Lee et al., 2024; BehnamGhader et al., 2024; Wang et al., 2024a), which enhances their generalization across diverse retrieval scenarios and query styles.
-This, in turn, improves their robustness against query complexity.
-Reranking models: sensitivity to query complexity.
-Figure 6: Attention heatmap of cross-encoder model (bge-reranker-m3): effect of increasing query conditions.
-When query conditions are few (1-3), the cross-encoder effectively attends to relevant conditions in the document, as the number of conditions increases, attention becomes more diluted.
-Reranking models compute relevance by jointly processing the query and document, primarily through: (1) cross-encoders, which perform token-level relevance comparison through cross-attention (e.g., bge-reranker-v2-m3), and (2) generative models that estimate relevance using LLM-based agents (e.g., bge-reranker-v2-gemma and FollowIR).
-Both architectures rely on deep query-document interaction, making them more sensitive to input complexity, such as changes in condition quantity and query format.
-Our experiments show that rerankers outperform retrieval models in ranking effectiveness for single-condition queries, suggesting their advantage in capturing fine-grained query-document relevance for short and simple queries.
-However, as query complexity increases, their performance declines more sharply, eventually falling behind retrieval models.
-Table 1 shows that rerankers’ average Win Rate drops by 39.05%, more than twice the decline of retrieval models (14.06%).
-A case study in Fig.
-6 provides further insight into this decline by analyzing the attention heatmap of a cross-encoder reranker.
-When the query contains only a few conditions (e.g., 1–3), the model effectively aligns each condition with relevant document sentences.
-However, as conditions increase, attention becomes unevenly distributed, with earlier conditions receiving more focus while later ones are overlooked.
-This imbalance likely contributes to the sharp drop in Win Rate observed in Task 2—as later conditions receive insufficient attention, the model struggles to distinguish relevance variations in corresponding document sentences.
-These findings highlight a key limitation in existing reranking models: their reliance on fine-grained query-document interactions becomes a bottleneck in multi-condition scenarios.
-While reranking models outperform retrieval models for queries with fewer conditions, their effectiveness diminishes as query complexity grows.
-5.2 The Effect of Document Side
-Previous experiments focused on query-side variations. We now investigate how changes on the document side, particularly document length, affect retrieval effectiveness.
-Figure 5 shows the impact of document length (padded to 512 and 1024 words) on relevance monotonicity [^6] The complete results and the padding filler are provided in Appendix D.3.
-Compared to original documents (Fig.4), increasing document length does not change the overall trend in Task 2—as hard negatives (HNs) become more challenging, models struggle more to distinguish relevance between documents satisfying $k$ vs.
-$k-1$ conditions.
-This decline is more pronounced when document length increases to 1024 words.
-Rerankers are particularly affected, showing a steeper win rate drop than retrieval models.
-Since they process queries and documents jointly, longer inputs likely increase attention dispersion in cross-encoders and complexity in generative models, weakening relevance estimation.
-Among retrieval models, GritLM and Nv-embed demonstrate strong robustness to document length variations.
-At 512 words, GritLM (66.66%) and Nv-embed (62.91%) achieve the highest average win rate.
-When document length is increased to 1024 words, Nv-embed (63.41%) outperforms GritLM (59.46%), indicating its ability to retain essential information in longer documents.
-This suggests that Nv-embed is more effective at capturing key content even when document length increases.
-5.3 Robustness of GritLM and NV-Embed
+Our experiments reveal notable differences between retrievers and rerankers across the three tasks.
+Fine-tuned rerankers exhibited excellent ranking performance under single-condition queries, but their efficacy rapidly diminished as the number of conditions increased.
+Retrievers demonstrate greater robustness under multi-condition queries and query style variations.
+We hypothesize that one contributing factor to these performance disparities lies in the training datasets.
+Many dense retrieval models are trained on a mixture of retrieval-specific and general textual datasets (Lee et al., 2024; BehnamGhader et al., 2024; Wang et al., 2024a).
+Such diverse training enhances their generalization across various retrieval scenarios and query styles, which, in turn, improves their robustness against query complexity.
+Beyond training data, we posit that the distinct input processing mechanisms also contribute to the observed performance differences.
+Retriever models typically employ a bi-encoder architecture, processing queries and documents independently.
+Conversely, rerankers, which process a concatenation of the query and document as a single input, appear more susceptible to input complexification—whether arising from an increase in query conditions or changes in query style.
+To validate this speculation, we re-evaluated the Win Rate for the relevance monotonicity task under multi-condition queries using documents that padded to 512 and 1024 words.
+Results as shown in Fig.5 , revealed that fine-tuned rerankers are highly sensitive to such increases in document length, which further illustrates the sensitivity of rerankers to complex input.
+In contrast, retriever models demonstrated greater robustness to length modifications.
+5.2 Condition Position Impact on Focus
 (a) Mean pooling
@@
 The dashed line represents the original data, while the solid line shows the Gaussian-smoothed trend (kernel size = 1) for clarity.
-GritLM and hybrid attention.
-GritLM exhibits strong robustness in multi-condition retrieval, which we hypothesize is due to its hybrid attention mechanism, integrating causal attention for generative task and bidirectional attention for embedding learning.
-Recent work has incorporated bidirectional attention into LLM-based retrievers, such as LLM2Vec (BehnamGhader et al., 2024) and Nv-Embed (Lee et al., 2024), which discard causal attention after converting LLMs into embedding models.
-In contrast, GritLM preserves causal attention, maintaining generative reasoning capabilities alongside representation learning.
-This likely explains its superior adaptability in multi-condition retrieval.
-Empirical results further validate this advantage: GritLM achieves state-of-the-art performance in complex instruction-following retrieval (Zhou et al., 2025) and ranks second in reasoning-intensive retrieval tasks (Su et al., 2024), reinforcing the importance of hybrid attention in embedding-based retrieval.
-Nv-Embed and latent attention layer pooling.
-Nv-Embed performs suboptimally in most multi-condition retrieval tasks, which we hypothesize is due to its Latent Attention Layer Pooling strategy.
-Common embedding methods include: 1) mean pooling, and 2) the last <EOS> token embedding.
-Mean pooling averages all token representations to generate a global representation, but potentially dilute critical information.
-Last <EOS> token embedding relies solely on the final token, which could introduce recency bias.
-A case study in Fig.
-7 examines how different pooling methods handle condition placement in queries.
-The results show that mean pooling leads to decreasing similarity scores as the condition moves further back, while <EOS> pooling exhibits an increasing trend, confirming its recency bias.
-In contrast, latent attention layer pooling maintains relatively stable performance across different condition positions, indicating improved information retention.
-We validated the robustness of Nv-Embed in handling longer documents in Section 5.2.
-Notably, at 1024 words, it maintains consistent monotonic relevance, preserving key information more effectively than other models.
-These findings suggest that its latent attention layer pooling potentially mitigates information loss, enhancing retrieval stability in multi-condition scenarios.
+Our experimental findings indicate that the position of a condition significantly influences the model’s subsequent similarity judgment.
+This phenomenon was observed consistently across both retriever and reranker models.
+To illustrate position effects in retriever models, we conducted a targeted study on how relocating a single condition within the query influences its similarity score.
+We selected representative retriever models employing distinct pooling strategies: mean pooling, <EOS> pooling, and latent layer pooling.
+The results, as depicted in Fig.6, revealed that models utilizing mean pooling tend to weight the early tokens most heavily: similarity drops steadily as the condition is shifted toward the tail of the query.
+<EOS> shows the opposite bias, emphasising the final tokens.
+Latent layer pooling heightened focus on both the beginning and end of the query, with comparatively less focus on the middle.
+Figure 7: Attention heatmap of cross-encoder model (bge-reranker-m3).
+Similarly, for reranker models, we selected a cross-encoder model (bge-reranker-m3) to visualize the attention heatmap.
+Fig.7 shows a non-uniform distribution of attention across different token positions.
+This implies that these rerankers tend to assign differential focus to specific conditions or tokens within the concatenated query-document input, rather than distributing their attention uniformly across all elements.
 6 Conclusion
-In this paper, we introduce MultiConIR, the first benchmark specifically designed to evaluate retrieval models in multi-condition scenarios.
-Unlike existing datasets that primarily focus on single-condition queries from search engines, MultiConIR captures real-world complexity by incorporating five diverse domains: books, movies, people, medical cases, and legal documents.
-We propose three tasks to systematically assess retrieval and reranking models on multi-condition robustness, monotonic relevance ranking, and query format sensitivity.
-Through extensive experiments, we analyze model performance across architectures and sizes, examining their ability to distinguish fine-grained conditions, preserve semantic information, and adapt to contextual variations.
-Our findings reveal that existing retrieval and reranking models struggle with multi-condition retrieval, with rerankers suffering severe performance degradation as query complexity increases.
-We further investigate the performance gap between retrieval and reranking models, exploring potential reasons for these discrepancies.
-Finally, we highlight the strengths of GritLM and Nv-Embed, which demonstrate enhanced adaptability to multi-condition queries, offering insights for future retrieval models.
+In this work, we introduced MultiConIR, a novel benchmark designed to rigorously evaluate information retrieval models in realistic multi-condition scenarios, a critical area where existing evaluation frameworks are lacking.
+Through three specifically designed tasks—complexity robustness, relevance monotonicity, and query format sensitivity—conducted across five diverse domains.
+Experiments revealed that existing models struggle with multi-condition retrieval, with their performance degrading as the number of conditions increases; rerankers excel for single-condition queries but fail in multi-condition scenarios.
+Notably, rerankers are more sensitive to complex inputs.
+GPT-4o outperforms specialised IR systems, exposing a performance gap in handling complex information needs.
+Our findings highlight an urgent need for new modeling approaches and training paradigms specifically tailored for robust multi-condition understanding.
+MultiConIR serves as a valuable resource to drive this research, benchmark progress, and ultimately propel information retrieval systems towards a more sophisticated, human-like comprehension of complex information needs.
 Limitations
@@
 For each model used in this paper, Table 4 provides details on model size, architecture, maximum input context length, and whether instructions is included.
+The GPT-4o model utilized for dataset generation and ranking in this work was the "2024-07-01-preview" version.
 Model Size Architecture Instruction Max length Pooling Method
@@
 LLM2Vec (BehnamGhader et al., 2024) 7.5B Decoder Yes 8K Mean
-Reranker
+Fine-tuned Reranker
 bge-reranker-v2-m3 (Chen et al., 2024) 568M Cross-Encoder No 8k N/A
-bge-reranker-v2-gemma (Chen et al., 2024) 2.5B Decoder Yes 1k N/A
+bge-reranker-v2-gemma (Chen et al., 2024) 2.5B Decoder Yes 8K N/A
 followIR (Weller et al., 2024a) 7.2B Decoder Yes 4K N/A
+RankZephyr (Pradeep et al., 2023) 7B Decoder Yes 32K N/A
+Zero-shot LLM for Ranking
+GPT-4o (OpenAI, 2024) N/A Decoder Yes 128K N/A
 Table 4: Details of models used in experiments.
@@
 For models that do not require instructions, we directly input the query and document, such as jina-embeddings-v3, gte-large-en-v1.5, and bge-reranker-v2-m3.
-Appendix B Prompt Templates For Constructing MultiCon Dataset
-Table 4, 5, and 6 present the prompts used in Steps 1 to 3 for constructing our MultiCon dataset.
-[…]
+Appendix B Prompt Templates For Constructing Multiconir Dataset
+Table 5, 6, and 7 present the prompts used in Steps 1 to 3 for constructing our Multiconir dataset.
+For placeholders, {domain} $\in$ {People, Books, Movies, Medical Case, Legal Document}.
+{domain_features} specifies key attributes within a particular domain.
+In the medical case domain, {domain_features} $\in$ { patient symptoms, clinical diagnosis, drug allergies, family medical history, surgical details, postoperative outcomes, hospitalization duration, recovery status.} In the legal document domain, {domain_features} $\in$ { case type, involved parties, court ruling, legal provisions, evidence summary, defense strategy.
+} In the movies domain, {domain_features} $\in$ { summary, lead actors, release date, release area, genre, detailed plots.
+} In the books domain, {domain_features} $\in$ { author, publication year, genre, main content, detailed plots.
+} In the people domain, {domain_features} $\in$ { profession, nationality, notable achievements, social impact, related events.
+}
 Task Prompt
@@
 Table 7: Prompt for GPT-4o to modify the condition sentence to hard negative sentence (Step 3).
-Appendix C Details And Examples Of The MultiCon Dataset
-Table 7 presents the five domains and their source data used for constructing our MultiCon dataset. Tables 9, 10, 8, 11, and 12 illustrate examples from their respective domains.
-Domain Records Source Data
-People 420 People Wikipedia Dat(Mahajan, 2017)
-Books 482 Books Dataset (Rustamov, 2021)
-Movies 500 Wikipedia Movie Plots (Robischon, 2018)
-Medical Case 479 Medical Cases (HPE AI Solutions, 2023)
-Legal Document 426 LexGLUE (Chalkidis et al., 2022)
-Table 7: Domain-specific Source Data References. The Records column indicates the number of data entries for each domain in the MultiCon dataset.
-Query 10 Positive HN 1
+Appendix C Benchmark Quality Evaluation
+To guarantee the reliability of MultiConIR, we applied a two–stage audit that examines both query validity and document–label validity. Figure 3 gives a visual outline; full numbers appear in Table 1.
+Query validity.
+From each of the five domains (People, Books, Movies, Medical, Legal) we randomly sampled 100 multi-condition queries, yielding a 500-item evaluation set.
+Ten trained graduate annotators independently rated every query for linguistic naturalness, precision of constraints, and contextual plausibility (i.e., whether logically plausible).
+Inter-annotator agreement reached $93.7\%$ with Fleiss’ $\kappa=0.84$, indicating near-perfect consensus that the automatically generated queries resemble genuine information needs.
+Document–label validity.
+  1. 1.
+    LLM filtering.
+We applied GPT-4o to the entire corpus.
+For every positive document $d^{+}$ the model verified that all ten conditions in $C(d)$ were satisfied; for each hard-negative document $d^{-}_{k}$ it checked that exactly k1k\!-\!1 conditions held.
+Instances failing these criteria (false positives or false negatives) were discarded, reducing domain sizes to: People (420), Books (482), Movies (500), Medical (479), Legal (426).
+  2. 2.
+    Human spot-check.
+To confirm the LLM filter, we randomly drew another 100 document–query pairs per domain (500 in total).
+Two independent annotators judged whether the labelled number of satisfied conditions was correct; disagreements were resolved by a third adjudicator.
+The residual error rate was $2.4\%$, implying that the automatic filter removed the vast majority of mis-labelled items.
+Appendix D Discussion: Win Rate vs. Traditional IR Metrics
+Why introduce Win Rate?
+MultiConIR poses multi-condition queries for which models must sense fine-grained semantic differences.
+We focus on two abilities: (i) discriminating the positive document from a hard negative as the number of query conditions grows (Task 1); (ii) preserving a monotonic ordering in which a document satisfying $k$ conditions outranks one satisfying k1k\!-\!1 under the same query (Task 2).
+Limitations of conventional metrics.
+  • Precision@1 coincides with Win Rate in Task 1 (one positive vs.
+one HN) but, in Task 2, observes only the top result and ignores the intended hierarchy $d^{+}\succ\mathrm{HN}_{1}\succ\cdots\succ\mathrm{HN}_{10}$.
+  • NDCG@$k$ introduces graded relevance but still weights absolute rank more than pairwise consistency, thus blurring step-wise violations of the monotonic order.
+  • Recall proved even less informative in early pilot runs dominated by easy negatives: high recall was achievable without respecting the semantic precision that MultiConIR is designed to test.
+How Win Rate fills the gap.
+Win Rate computes the proportion of pairwise comparisons in which a document that fulfils more conditions is ranked above one that fulfils fewer.
+Hence, it matches Precision@1 in the degenerate Task 1 case, yet remains sensitive to every local inversion in the graded Task 2 ladder, offering a sharper lens on a model’s ability to capture incremental semantic constraints.
+Appendix E Examples Of The MultiConIR Dataset
+Tables 9, 10, 8, 11, and 12 illustrate examples from their respective domains.
+Query 3 Positive HN 1
 Find a notable individual who meets these criteria: 1.
@@
 Worked at NIMH in Bethesda, Maryland before 1981.
-4.
-Collaborated closely with Robert Sack in Oregon.
-5.
-Researches chronobiologic sleep and mood disorders.
-6.
-Graduated from University of Chicago in nineteen seventies with MD and PhD.
-7.
-Had around 90 publications on PubMed by 2005.
-8.
-Is a full professor and vice-chair of Psychiatry at OHSU.
-9.
-Focuses on bright light exposure and melatonin treatments.
-10.
-Studies disorders like winter depression, jet lag, and shift work maladaptation.
 Relying on a very precise assay for plasma melatonin, a hormone that has a clearly defined 24-hour pattern of secretion, biological rhythm disorders can be assessed and their treatment can be monitored.
@@
 Totally blind individuals have 25-hour circadian rhythms, drifting an hour later each day unless they take a melatonin capsule at a certain time every day.
-Prior to moving to Oregon in 1981, Lewy was at the National Institute of Mental Health (NIMH) in Bethesda, Maryland, working with senior colleague Thomas Wehr.
+After to moving to Oregon in 1981, Lewy was at the National Institute of Mental Health (NIMH) in Bethesda, Maryland, working with senior colleague Thomas Wehr.
 In Oregon, he has worked closely with Robert L.
@@
 Current research is focused on developing bright light exposure and melatonin administration as treatment modalities for these disorders.
-These lifestyle changes address issues like winter depression, jet lag, shift work maladaptation, and certain types of sleep disturbances.
+These disorders include winter depression, jet lag, maladaptation to shift work, and certain types of sleep disturbances.
 Table 8: An example in domain of People
-Query 10 Positive HN 1
+Query 5 Positive HN 1
 Find a notable individual who meets these criteria: 1.
@@
 Contains exclusive news and material.
-6.
-Includes Bernie’s Little Black Book.
-7.
-Features Madoff’s past calendars.
-8.
-Provides details on the origins of his scam.
-9.
-Includes cooperation from Eleanor Squillari.
-10.
-Reveals SEC connections to Madoff.
 The collapse of Bernie Madoff’s Ponzi scheme led to the instant evaporation of \$65 billion of wealth.
@@
 This is an incisive and voyeuristic look into this first family of financial crime.
-The Madoff Chronicles includes a vast array of news and material that readers won’t find anywhere else.
+The Madoff Chronicles includes a vast array of news and material.
 Contains a reproduction of Bernie’s Little Black Book.
@@
 Extensive cooperation by Madoff’s personal assistant, Eleanor Squillari.
-Contains no documented communications between Madoff and SEC members.
+Contains incriminating connections between Madoff and certain members of the SEC.
 Table 9: An example in domain of Book
-Query 10 Positive HN 1
+Query 8 Positive HN 1
 Find a movie that matches all conditions: 1.
@@
 Plot: Terry’s spending takes a toll.
-9.
-Plot: Lottie sees Terry and Julia at park.
-10.
-Release Year: 1918.
 Origin/Ethnicity: American
@@
 Table 10: An example in domain of Movie
-Query 10 Positive HN 1
+Query 7 Positive HN 1
 Find a case where the patient: 1.
@@
 Experienced TIA and moderate carotid stenosis.
-8.
-Had 400 ml blood loss.
-9.
-Provided informed consent for the procedure.
-10.
-Received a 6-French sheath in the right femoral artery.
 PROCEDURE PERFORMED:
@@
 Bilateral cerebral angiograms were performed as well via right and left common carotid artery injections.
-INDICATIONS FOR PROCEDURE: TIA, aortic stenosis, postoperative procedure. Moderate carotid artery stenosis.
+INDICATIONS FOR PROCEDURE: TIA, aortic stenosis, postoperative procedure. Severe carotid artery stenosis.
 ESTIMATED BLOOD LOSS: 400 ml.
@@
 Table 12: An example in domain of Legal Document
-Appendix D Complete Results
-D.1 Complete Results Of Task1
-Table 13, 14, 15, 16, and 17 present the complete results of Task 1 for the five datasets: People, Books, Movies, Medical Case, and Legal Document.
-Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 Decline
-Spare Retriever
-BM25 23.10 25.24 22.62 26.43 26.19 23.57 24.05 22.14 22.38 23.81 -0.71
-Dense Retriever
-jina-embeddings-v3 88.81 78.33 78.33 76.90 72.14 68.57 67.38 66.19 67.62 65.95 22.86
-gte-large-en-v1.5 72.14 76.19 73.10 68.57 69.29 61.43 66.19 62.38 69.05 63.81 8.33
-NV-Embed-v2 82.14 89.52 85.48 85.48 87.38 84.05 87.38 84.05 84.05 86.67 -4.52
-bge-en-icl 85.00 85.95 86.43 85.00 83.81 84.29 86.43 87.14 85.48 83.81 1.19
-gte-Qwen2-7B-instruct 58.10 70.95 69.05 70.71 71.90 71.43 71.43 69.76 72.62 69.76 -11.67
-gte-Qwen2-1.5B-instruct 61.90 74.76 70.24 71.19 73.57 70.48 73.81 66.90 65.71 65.00 -3.10
-e5-mistral-7b-instruct 76.43 77.86 73.33 75.48 68.10 71.43 67.62 68.81 69.05 73.57 2.86
-GritLM-7B 79.05 84.52 80.71 81.67 81.67 77.86 83.10 82.62 81.43 82.62 -3.57
-LLM2Vec 79.05 74.76 71.90 75.71 74.05 75.95 75.00 78.81 76.19 80.00 -0.95
-Reranker
-bge-reranker-v2-m3 90.00 85.95 78.57 73.81 74.29 62.14 63.57 56.67 52.14 36.67 53.33
-bge-reranker-v2-gemma 92.62 93.33 91.67 88.33 89.29 84.76 81.43 72.38 69.52 50.71 41.90
-followIR 84.52 80.48 76.43 68.57 60.71 56.90 50.00 34.05 20.95 8.57 75.95
-Table 13: Win Rate Score as the Number of Conditions in the Query Increases on the People Dataset (Task 1)
-Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 decline
-Spare Retriever
-BM25 45.85 47.93 47.93 49.17 47.72 51.24 50.21 48.34 53.94 46.06 -0.21
-Dense Retriever
-jina-embeddings-v3 76.14 72.41 71.99 71.37 67.01 64.73 67.43 63.07 63.28 58.51 17.63
-gte-large-en-v1.5 73.86 75.93 71.99 70.33 68.67 69.71 65.77 66.60 64.32 70.95 2.90
-NV-Embed-v2 78.01 73.65 75.93 68.88 69.71 69.50 68.46 68.05 67.84 71.58 6.43
-bge-en-icl 78.42 76.76 75.31 72.20 72.41 69.29 70.54 70.12 70.95 72.41 6.02
-gte-Qwen2-7B-instruct 71.99 63.90 64.52 61.83 54.77 50.00 50.00 49.17 48.55 50.83 21.16
-gte-Qwen2-1.5B-instruct 70.54 69.09 70.33 69.92 65.98 60.58 64.11 62.45 64.11 56.85 13.69
-e5-mistral-7b-instruct 73.03 65.77 68.88 65.98 60.58 59.96 59.34 60.58 60.37 68.67 4.36
-GritLM-7B 80.91 74.69 81.33 73.86 75.73 73.86 79.25 77.80 78.01 80.50 0.41
-LLM2Vec 80.08 70.12 72.41 70.12 68.05 66.80 64.32 65.35 65.15 64.52 15.56
-Reranker
-bge-reranker-v2-m3 83.20 79.05 76.35 71.99 72.20 72.82 74.48 68.67 60.58 51.66 31.54
-bge-reranker-v2-gemma 87.76 84.44 82.78 81.74 80.08 75.31 74.48 72.82 70.75 70.54 17.22
-followIR 85.89 78.22 77.59 72.41 68.67 68.46 61.41 60.37 53.11 47.30 38.59
-Table 14: Win Rate Score as the Number of Conditions in the Query Increases on the Books Dataset (Task 1)
-Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 decline
-Spare Retriever
-BM25 28.60 34.60 33.60 37.20 34.80 36.80 37.00 38.40 37.60 39.00 -10.40
-Dense Retriever
-jina-embeddings-v3 77.60 77.00 75.20 75.40 70.80 72.00 69.40 68.40 71.80 67.20 10.40
-gte-large-en-v1.5 78.40 81.00 76.60 75.60 73.80 71.00 71.00 69.00 67.20 59.40 19.00
-NV-Embed-v2 81.00 81.80 80.40 78.60 80.00 76.00 77.40 76.60 72.20 67.00 14.00
-bge-en-icl 84.40 83.20 80.60 79.20 78.80 77.00 76.00 75.80 70.40 66.60 17.80
-gte-Qwen2-7B-instruct 74.60 80.00 74.60 72.80 72.40 72.00 68.80 75.00 67.60 67.40 7.20
-gte-Qwen2-1.5B-instruct 80.40 83.00 76.00 74.80 73.00 73.80 71.00 74.60 72.20 68.40 12.00
-e5-mistral-7b-instruct 73.40 71.80 66.00 64.80 70.00 61.20 65.60 62.80 63.00 55.60 17.80
-GritLM-7B 85.00 83.60 80.00 80.80 79.80 72.80 76.40 72.60 71.40 73.80 11.20
-LLM2Vec 87.20 89.40 86.40 84.40 81.80 81.60 79.60 77.80 75.40 74.20 13.00
-Reranker
-bge-reranker-v2-m3 92.80 90.80 80.20 81.60 79.20 75.60 71.00 63.40 67.00 60.40 32.40
-bge-reranker-v2-gemma 92.20 91.60 86.00 89.20 83.40 79.40 77.60 78.20 70.80 61.60 30.60
-followIR 92.60 86.40 83.40 88.00 81.00 81.00 83.40 81.40 82.60 85.00 7.60
-Table 15: Win Rate Score as the Number of Conditions in the Query Increases on the Movies Dataset (Task 1)
-Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 Decline
-Spare Retriever
-BM25 25.68 28.18 31.52 32.99 37.37 36.33 35.49 39.25 37.16 40.71 -15.03
-Dense Retriever
-jina-embeddings-v3 68.89 64.93 65.34 65.55 55.32 59.08 56.16 63.67 60.54 59.29 9.60
-gte-large-en-v1.5 77.24 81.84 76.83 75.57 71.82 70.56 70.35 67.01 68.48 62.42 14.82
-NV-Embed-v2 77.24 73.07 74.53 73.70 65.97 62.63 67.43 68.68 67.22 63.67 13.57
-bge-en-icl 80.79 78.91 76.41 77.04 68.89 68.68 72.23 71.19 69.31 67.64 13.15
-gte-Qwen2-7B-instruct 71.61 73.70 73.49 72.65 66.60 67.64 66.18 64.72 60.75 51.57 20.04
-gte-Qwen2-1.5B-instruct 73.90 72.86 72.03 70.77 67.43 65.55 63.47 64.09 60.13 53.86 20.04
-e5-mistral-7b-instruct 69.31 67.01 67.64 64.09 57.83 55.53 56.16 51.98 50.10 48.64 20.67
-GritLM-7B 79.75 82.25 76.62 78.08 77.66 75.16 75.16 77.87 78.91 78.29 1.46
-LLM2Vec 84.34 81.21 79.12 83.09 80.58 79.75 78.71 79.33 80.79 78.50 5.85
-Reranker
-bge-reranker-v2-m3 81.21 91.02 86.64 86.43 79.33 75.37 75.78 66.39 72.65 62.00 19.21
-bge-reranker-v2-gemma 89.35 92.48 90.19 92.90 91.02 85.80 85.39 82.67 80.79 70.35 19.00
-followIR 92.28 90.81 89.35 89.56 86.22 84.97 81.42 86.22 76.41 71.82 20.46
-Table 16: Win Rate Score as the Number of Conditions in the Query Increases on the Medical Case Dataset (Task 1)
-Model Query1 Query2 Query3 Query4 Query5 Query6 Query7 Query8 Query9 Query10 decline
-Spare Retriever
-BM25 19.72 34.04 30.05 36.85 40.85 41.78 43.66 43.90 43.19 49.77 -30.05
-Dense Retriever
-jina-embeddings-v3 69.01 63.62 68.31 65.96 62.68 63.85 63.38 59.86 59.86 52.58 16.43
-gte-large-en-v1.5 77.70 71.36 70.42 61.03 69.95 64.32 64.32 59.86 57.75 50.23 27.46
-NV-Embed-v2 84.27 83.57 77.70 71.83 75.35 70.89 65.73 60.33 58.69 51.17 33.10
-bge-en-icl 88.50 78.40 73.47 70.42 68.78 65.73 65.96 61.97 52.35 49.53 38.97
-gte-Qwen2-7B-instruct 77.46 72.54 68.31 64.55 60.33 56.57 54.69 52.35 46.24 41.31 36.15
-gte-Qwen2-1.5B-instruct 81.46 75.12 72.54 70.19 69.72 66.90 62.21 55.16 56.10 49.30 32.16
-e5-mistral-7b-instruct 83.10 71.83 65.02 66.90 61.97 59.86 49.77 51.17 46.71 44.13 38.97
-GritLM-7B 85.68 76.53 73.24 67.61 67.14 67.84 64.55 62.21 62.91 64.55 21.13
-LLM2Vec 84.98 71.60 67.61 64.08 57.98 58.69 55.16 52.35 46.01 37.79 47.18
-Reranker
-bge-reranker-v2-m3 88.50 80.99 71.36 66.43 66.43 56.10 54.46 42.25 25.59 13.62 74.88
-bge-reranker-v2-gemma 93.43 88.26 82.86 72.77 72.07 69.72 60.56 55.40 43.66 27.23 66.20
-followIR 61.74 62.68 54.46 54.46 53.99 48.36 36.85 17.61 9.86 4.93 56.81
-Table 17: Win Rate Score as the Number of Conditions in the Query Increases on the Legal Documents Dataset (Task 1)
-D.2 Complete Results Of Task 2
+Appendix F Complete Results
+F.1 Complete Results Of Task 2
 Table 13 presents the experimental results of Task 2, where Win Rate reflects the success rate between documents that satisfy different numbers of conditions under a multi-condition query (query10, which contains ten conditions), i.e., $d_{k}$ vs.
 $d_{k-1}$.
-Model $d_{1}$_vs_$d_{0}$ $d_{2}$_vs_$d_{1}$ $d_{3}$_vs_$d_{2}$ $d_{4}$_vs_$d_{3}$ $d_{5}$_vs_$d_{4}$ $d_{6}$_vs_$d_{5}$ $d_{7}$_vs_$d_{6}$ $d_{8}$_vs_$d_{7}$ $d_{9}$_vs_$d_{8}$ $d_{10}$_vs_$d_{9}$ Avg.
-Spare Retriever
+[…]
+Sparse Retriever
 BM25 13.91 16.50 16.81 18.14 22.10 29.04 37.78 38.87 39.93 40.19 25.90
@@
 LLM2Vec 83.50 74.25 73.43 72.24 70.36 67.21 66.99 67.07 66.49 67.48 70.90
-Reranker
+Fine-tuned Reranker
 bge-reranker-v2-m3 76.08 68.06 63.83 62.06 60.65 58.35 54.79 50.60 48.57 44.96 58.80
@@
 followIR 61.99 59.82 60.87 60.76 59.91 56.41 51.63 47.93 44.71 43.52 54.76
+RankZephyr 90.20 86.04 83.96 83.14 82.23 79.41 76.07 72.43 70.26 66.84 79.06
+Zero-shot LLM for Ranking
+GPT-4o 93.37 92.76 92.20 91.16 90.51 89.55 88.38 86.82 85.96 85.26 89.60
 Table 13: Average Win Rate Comparison Between Documents in Task 2
-D.3 Complete Results Of Document Length
+F.2 Complete Results Of Document Length
 Table 14 and Table 15 present the effect of document length on retrieval performance, with documents padded to 512 and 1024 words, respectively.
@@
 The filler text is inserted between the original document sentences until the total text length reaches 512 or 1024 words.
-Model $d_{1}$_vs_$d_{0}$ $d_{2}$_vs_$d_{1}$ $d_{3}$_vs_$d_{2}$ $d_{4}$_vs_$d_{3}$ $d_{5}$_vs_$d_{4}$ $d_{6}$_vs_$d_{5}$ $d_{7}$_vs_$d_{6}$ $d_{8}$_vs_$d_{7}$ $d_{9}$_vs_$d_{8}$ $d^{+}$_vs_$d_{9}$ Avg.
-Spare Retriever
+[…]
+Sparse Retriever
 BM25 11.91 14.34 14.75 14.45 15.99 24.75 36.58 37.68 38.02 39.34 24.78
@@
 LLM2Vec 70.57 68.37 67.10 67.10 66.35 58.73 36.04 36.81 35.37 37.03 54.35
-Reranker
+Fine-tuned Reranker
 bge-reranker-v2-m3 73.65 66.55 63.28 61.65 54.60 38.54 28.42 27.92 28.14 28.02 47.08
@@
 followIR 50.48 51.78 49.64 46.27 37.60 25.22 24.41 24.51 25.67 24.71 36.03
+RankZephyr 85.41 79.91 72.91 70.41 64.41 52.91 47.91 45.41 42.30 42.56 60.41
+Zero-shot LLM for Ranking
+GPT-4o 89.41 88.15 87.50 87.80 86.30 84.90 83.20 82.95 81.10 80.35 85.17
 Table 14: Effect of document length on retrieval performance (padded to 512 words).
-Model $d_{1}$_vs_$d_{0}$ $d_{2}$_vs_$d_{1}$ $d_{3}$_vs_$d_{2}$ $d_{4}$_vs_$d_{3}$ $d_{5}$_vs_$d_{4}$ $d_{6}$_vs_$d_{5}$ $d_{7}$_vs_$d_{6}$ $d_{8}$_vs_$d_{7}$ $d_{9}$_vs_$d_{8}$ $d^{+}$_vs_$d_{9}$ Avg.
-Spare Retriever
+[…]
+Sparse Retriever
 BM25 12.25 14.32 14.86 14.81 15.97 24.48 36.00 37.53 38.85 39.53 24.86
@@
 LLM2Vec 74.81 72.05 71.55 26.85 26.68 26.35 25.24 26.16 26.04 27.27 40.30
-Reranker
+Fine-tuned Reranker
 bge-reranker-v2-m3 76.56 70.75 56.83 18.33 19.86 19.09 18.55 20.75 19.88 21.78 34.24
@@
 followIR 52.36 51.43 19.70 18.35 17.15 17.11 17.78 16.94 17.77 18.75 24.73
+RankZephyr 84.34 76.84 66.64 35.14 33.34 34.04 31.54 32.64 30.74 31.24 45.65
+Zero-shot LLM for Ranking
+GPT-4o 88.80 86.50 85.80 86.10 83.50 82.00 80.20 79.50 78.30 77.63 82.83
 Table 15: Effect of document length on retrieval performance (padded to 1024 words).
-Appendix E Findings In Constructing MultiCon Dataset
-E.1 The Use Of LLM-generated Data in Retrieval
-In recent years, artificial datasets generated by LLMs have become a common practice for training and evaluating retrieval models (Su et al., 2024; Lee et al., 2024; Weller et al., 2024a).
+F.3 Comparative Results for other LLMs
+We evaluated two additional LLMs—Qwen3-32B and Gemini-2.0-Flash—on the same MultiConIR Task 1.
+As shown in Table 16, the results indicate that Gemini-2.0-Flash attains the highest WinRate (91.20), marginally outperforming GPT-4o (90.47), with Qwen3-32B in third place (86.61).
+GPT-4o is therefore not significantly superior, confirming that MultiConIR does not inherently favor GPT-4o.
+Model Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Q9 Q10 Avg.
+Qwen3-32B 91.20 90.64 89.55 88.08 86.91 85.35 84.74 84.44 83.27 81.88 86.61
+GPT-4o 95.49 94.89 93.71 92.11 90.81 89.08 88.43 88.08 86.82 85.26 90.47
+Gemini-2.0-flash 95.67 95.10 94.06 92.63 91.50 89.99 89.38 89.08 87.95 86.61 91.20
+Table 16: Performance comparison across Q1–Q10 queries and average score.
+Appendix G Findings In Constructing Multiconir Dataset
+G.1 The Use Of LLM-generated Data in Retrieval
+In recent years, artificial datasets generated by LLMs have become a common practice for training and evaluating retrieval models (Weller et al., 2024a; Shi et al., 2025, 2024; Zhang et al., 2025; Han et al., 2025).
 For instance, E5-Mistral (Wang et al., 2024a) rely entirely on LLM-generated datasets for fine-tuning.
 While this approach can significantly expand training corpora, prior studies have highlighted its potential drawbacks, including introducing inherit linguistic biases of the underlying LLMs (Shumailov et al., 2024), potentially constraining the retrieval model’s performance and generalizability.
-Furthermore, purely artificial data often lacks the contextual richness and complexity found in real-world retrieval scenarios (Li et al., 2023c; Wang et al., 2024b), making it difficult to capture the actual needs of users’ queries accurately.
+Furthermore, purely artificial data often lacks the contextual richness and complexity found in real-world retrieval scenarios (Li et al., 2023c; Wang et al., 2024c), making it difficult to capture the actual needs of users’ queries accurately.
 During our dataset construction, we observed similar issues.
@@
 This targeted approach preserves real-world data authenticity while introducing controlled semantic perturbations, ensuring that retrieval models are trained on meaningful and realistic hard negatives rather than fully synthetic documents.
-E.2 Impact of Different Hard Negative Construction Strategies
+G.2 Impact of Different Hard Negative Construction Strategies
 To systematically examine the impact of hard negative sentence (HNS) construction on retrieval models, we experimented with two distinct approaches: (1) Key Information Modification – altering critical details while maintaining overall sentence structure (applied to books, movies, medical cases, and legal documents).